### v2

# Explain What You Mean: Intent Augmented Knowledge Graph Recommender Built With An LLM

Abstract.
Interaction sparsity is a long-standing challenge in recommendation systems. Sparsity manifests in environments with disproportional cardinality of groupings of entities, such as users and products in an online marketplace. It is also found for newly introduced entities, described as the cold-start problem. Recent efforts to mitigate this issue either enrich the connectivity data by incorporating social networks or external knowledge graphs, or fine-tune LLMs into interaction augmenters or next-item recommenders. However, these techniques tend to be resource demanding, requiring high computational power. They also have several limitations, including data availability, low quality, or synthetic noise issues. In this work, we propose LLM-based Intent Knowledge Graph Recommender (IKGR), a novel framework that leverages retrieval-augmented generation and an encoding approach to construct and densify a knowledge graph. IKGR leverages latent user-item affinities from an interaction knowledge graph and further densifies it through mutual intent connectivity. This addresses sparsity issues and allows the model to make intent-grounded recommendations with an interpretable embedding translation layer. Through extensive experiments on real-world datasets, we demonstrate that IKGR overcomes knowledge gaps and achieves substantial gains over state-of-the-art baselines on both publicly available and our internal recommendation datasets.
Recommendation system, Large Language Model (LLM), Knowledge Graph, Retrieval Augmented Generation (RAG)
1. Introduction
Recommender systems fulfill a crucial role by helping users navigate through an ever-increasing catalog of online content and services to find what is most relevant for their personalized interests (Meng et al., 2023a; Wei et al., 2022). In enterprise applications, recommender algorithms are also widely deployed in internal search platforms or forums to help employees and developers ease the understanding of enterprise knowledge and the discovery of helpful articles, tips, or properties (Verma et al., 2015; Roy and Dutta, 2022).
A wide family of recommendation strategies has demonstrated success in modeling users’ historical interactions, such as collaborative filtering (CF) (Koren et al., 2021), graph-based methods (Pei et al., 2019; Liu et al., 2020; Zuo et al., 2024), and more recently recommenders based on Large Language Models (LLMs) (Kim et al., 2024; Bao et al., 2023; Wu et al., 2024). As these methods rely on implicit feedback from previous interactions, their performance tends to degrade for cold-starting new users or items with sparse interactions, hence it has always been a central topic in recommender systems to efficiently leverage user/item features and interactions to overcome sparse implicit feedback.
To address the cold start problem, researchers have proposed various approaches to augment the plain user-item interaction data with additional information. Depending on the source of added contextual information, these augmentation methods can be grouped into three categories: (1) enhancing user/item attributes with existing data, such as item attributes (Ying et al., 2023), user-generated content (Fan et al., 2019; Meng et al., 2023b), or multi-modal features (Yi et al., 2022); (2) enhancing linkages with existing data, such as social networks (Jamali and Ester, 2010) or external knowledge graphs (Wang et al., 2018), and (3) enhancing user/item links and attributes with LLMs, such as interaction augmentation (Wei et al., 2024). These approaches attempt to enrich entity attributes, increase interactions, or add contextual information to help the recommender learn better representations and provide improved recommendations. However, these augmentation methods can still face several challenges, which we categorize as follows.
i) Synthetic data noise caused by preference differences and knowledge gaps. Handling noisy data is a long-standing challenge faced by recommender systems trained on implicit feedback, where temporary or individualized user preferences can induce false positive or false negative samples. The noise issue can be further exacerbated by certain attempts to alleviate sparsity, such as the generation of synthetic user interactions with LLMs (Wei et al., 2024). LLM-generated preferences can introduce additional distribution skewness, potentially over-representing popular items, and failing to capture the long-tail distribution of real-world user preferences. To mitigate the additional data noise, careful calibration techniques need to be employed, e.g., loss calibration and training sample selection (Wei et al., 2024). Furthermore, LLMs may exhibit knowledge gaps, lacking necessary domain-specific context. In enterprise use cases especially, there is often a gap between private enterprise data and public corpora, leading to difficulty for both LLMs and traditional data augmentation approaches.
ii) Excessive resource budget. Another line of research has leveraged LLMs as an end-to-end function to perform the recommendation task (Wu et al., 2024; Deldjoo et al., 2024; Bao et al., 2023; Kim et al., 2024). However, fine-tuning the LLM to mitigate the knowledge gap tends to become resource intensive both in terms of the computation budget and developer efforts to calibrate the data and the model for each specific use case.
iii) Lack of fine-grained understanding. Among the LLM-based linking approaches, their predictions are mostly based on natural language representations rather than fact-based knowledge graph structures, hindering interpretability and targeted adjustments. For example, when recommending popular head items, it may not be clear if the recommendation is based on a genuine user affinity or simply the popularity of the item. A bias that is hard to correct without understanding the LLM’s underlying reasoning. Furthermore, as the number of real user interactions grows over time, the relative value of LLM-synthesized edges may diminish. Existing node enrichment methods use side information to augment node attributes with text-based features, but typically do not decompose them into graph structures, preventing the use of graph-based inference techniques (Ko et al., 2022; Liang et al., 2023).
Figure 1. In enterprise search, the queries could contain special terminologies and acronyms, where traditional search engines or personalized rerankers fail to capture the real intents under such knowledge gap. IKGR addresses the challenge via injecting fine-grained understanding to textual features.
Additionally, these approaches may generate abbreviated textual descriptions, which can impair the model’s representations and subsequent recommendations.
In order to build a recommender that is good at handling the knowledge gap mentioned above, this work proposes to leverage the general reasoning ability of the LLM to augment the interactions. We take the the enterprise search as a primary case study as visualized in Figure 1, while keeping our approach applicable to more general scenarios. Our objective is to address the issue of noisy LLM generations under knowledge gap while maximally reducing the computation budget. We propose Intent Knowledge Graph Recommender (IKGR), an LLM-based pipeline, to explicitly augment node-level information. We build a knowledge graph that is naturally linked to the user-item interaction graph, as a way to enhance user/item node side information. Inspired by their impressive natural language understanding ability, we utilize LLMs to augment the interaction graph. We also simplify the task for the LLM to only boost the node-level understanding, so that more reliable results can be expected. In the proposed IKGR pipeline, the LLM is asked to perform a node-level retrieval task, which returns a list of entities that reflect users’ consumption intents. To overcome the knowledge gap, we carefully prepare an external knowledge base, explaining common abbreviations and terms. To reduce the computation budget, we design the framework as tuning-free, while retrieving the knowledge explanations from the knowledge base and feed to the LLM via in-context learning. We also design the backend method to be a lightweight Graph Neural Network (GNN) based on embedding translation, which naturally leverages the learned intent representation to inject structure prior into the graph.
As the main contribution of this work, we see two major advantages of using an LLM to augment the knowledge graph in the IKGR pipeline. Firstly, IKGR embraces the shift from an heuristic embedding learning structure to an LLM-based explicit knowledge extraction regime. It leverages the LLM to predict a new type of node: the interaction intent from a natural language perspective. Unlike previous approaches that rely on uninterpretable vector embeddings to capture multifaceted node attributes, IKGR recognizes that sub-node intents written in natural language format are often extractable from node features based on LLM. The natural language representation provide a more intuitive reflection of user preferences and interaction intents compared to indirect embeddings, and could be more flexible in terms of intent quantities. By incorporating LLMs, IKGR captures the richness and context of natural language, enhancing the accuracy and effectiveness of recommendations. Secondly, the low-quality and incomplete side information is enhanced by leveraging the extensive knowledge of LLMs gained through the pre-training stage, which allows for a tuning-free model that is both aware of vast real-world knowledge and understands user preferences and interaction intents. This integration prevents the dispersion of features across separate vector spaces and provides more accurate results.
  • We tackle the data sparsity challenge in recommendation systems, and propose IKGR, a knowledge graph augmentation approach with the LLM. IKGR leverages in-context learning with LLM to mitigate knowledge gap, and focuses on the node-level intent understanding and linkage to avoid introducing noise in synthesizing interactions.
  • The proposed IKGR is LLM-tuning-free, and is able to build knowledge graph inherently linked to the user-item interactions without the need for entity alignment. IKGR further leverages the learned intent nodes to inject structural prior in the embedding space with a knowledge graph translation layer.
  • Results on the enterprise search dataset demonstrate that the proposed IKGR effectively mitigate the knowledge gap, and outperforms the state-of-the-art baselines on search engine optimization reranking. Results on open-sourced datasets also showed the superiority of IKGR on the general recommendation scenario.
2. Related Works
Personalized Re-ranking Recommender System. Re-ranking in recommender systems aims to refine an initially ranked list to better capture user preferences and item relationships. Traditional ranking methods optimize global performance but often overlook the mutual influence between items and user-specific intent (Pei et al., 2019). Recent approaches address these limitations by integrating semantic tag information (Liang et al., 2010), multi-stage ranking optimization (Evnine et al., 2024), and item relationships (Liu et al., 2020). Transformer-based models effectively model global item interactions, while graph-based methods leverage item relationships for improved ranking (Pei et al., 2019; Liu et al., 2020). Furthermore, self-supervised learning enhances sequential recommendation by mitigating data sparsity issues (Zhou et al., 2020). These methods collectively demonstrate that incorporating user personalization, item dependencies, and efficient ranking strategies significantly enhances re-ranking effectiveness.
Recommender Approaches Discovering Intent. Recent works explore intent modeling by leveraging external signals such as search queries and knowledge graphs. UDITSR (Zhang et al., 2024) jointly models search and recommendation by utilizing explicit search queries to infer implicit demand intents, while a dual-intent translation mechanism captures relationships between inherent intent, demand intent, and item interactions. Alternatively, knowledge graph-based methods, such as KGIN (Wang et al., 2021), refine intent discovery by representing user-item interactions through fine-grained relational paths, improving interpretability and recommendation quality. These approaches demonstrate that integrating external intent signals and structured relational modeling enhances intent-aware recommendations.
Recommender Approaches Addressing Data Sparsity. Recent works tackle the data sparsity issue through enhanced model architectures, graph-based techniques, and knowledge-aware methods. The Item History Model (IHM) improves cold-start item recommendations by directly injecting user-interaction data into the item tower and employing an inductive structure for real-time inference (Zuo et al., 2024). Cold Brew addresses sparsity in graph neural networks by distilling node representations, mitigating the impact of missing or noisy neighbors (Zheng et al., 2021). RippleNet further alleviates sparsity by propagating user preferences through knowledge graph relations, enriching item representations beyond collaborative signals (Wang et al., 2018). These approaches demonstrate that leveraging historical interactions, distilling structural knowledge, and integrating external information effectively mitigate data sparsity in recommendation systems.
LLM-Based Interaction Augmentation for Recommender Systems. Gen-RecSys provides a comprehensive overview of generative models in recommendation, highlighting LLM-driven natural language understanding and multimodal integration (Deldjoo et al., 2024). LLMRec introduces graph augmentation strategies using LLMs to enrich interaction graphs, refine side information, and denoise implicit feedback, demonstrating performance gains across benchmark datasets (Wei et al., 2024). BLAIR further bridges language and recommendation by pretraining sentence embeddings on large-scale review data, improving item retrieval in complex natural language contexts (Hou et al., 2024). Additionally, BERT4Rec employs bidirectional self-attention to model user behavior sequences, overcoming limitations of traditional sequential models (Sun et al., 2019). These studies collectively showcase the potential of LLMs in augmenting interactions, refining representations, and enhancing recommendation performance.
Directly Leveraging LLM as the Recommender. A-LLMRec enhances LLM-based recommendation by leveraging embeddings from state-of-the-art collaborative filtering models, excelling in both cold and warm scenarios while maintaining efficiency and model-agnostic integration (Kim et al., 2024). CoRAL introduces collaborative retrieval-augmented prompting, addressing LLMs’ reliance on semantic information by incorporating user-item interactions through reinforcement learning-based retrieval policies, significantly improving long-tail recommendation (Wu et al., 2024). TALLRec proposes a tuning framework to align LLMs with recommendation-specific tasks, demonstrating strong generalization and efficiency even with limited training data (Bao et al., 2023). These approaches collectively highlight the potential of LLMs as standalone recommenders by addressing cold-start challenges, enhancing collaborative reasoning, and improving alignment with recommendation-specific objectives.
3. Preliminaries
This work brings techniques from recommendation systems, knowledge graph-based GNNs, and Retrieval Augmented Generation (RAG) with LLMs. To help readers better digest our terminology and approaches, we leverage this section to present the problem formulation and commonly adopted techniques across these domains, and introduce our contributions in Section 4.
3.1. Graph Formulation Of Recommendation System
We consider the recommendation system where a set of registered users $\mathcal{U}$ interact over a set of items $\mathcal{I}$. Define the interactions as a graph $\mathcal{G}$, and each user $u$ or item $i$ is a node in the graph. The collection of user-item interactions are the edges in the graph, denoted as $\mathcal{E}$. We employ the implicit feedback protocol, where each edge of $(u,i)$ implies the user $u\in\mathcal{U}$ consumes the item $i\in\mathcal{I}$. The goal is to learn a model that recommends the top-N items for a target user.
The goal is to learn a scoring function that is trained over some graph $\mathcal{G^{\prime}}$, and predicts positive interactions among a set of $(u,i)$ pairs at inference time. This scoring function is denoted as
$g_{\Theta}(\{(u,i)\text{ pairs}\};\mathcal{G^{\prime}})$, parameterized by $\Theta$. In mathematical formulation, the objective of the recommendation systems is to maximize the link prediction posterior probability of accurately predicting all interactions in the dev set:
(1) $\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}})=\mathcal{E}_{\text{dev}})$
where $\{(u,i)\}_{\text{dev}}$ is the dev set input user-item pairs, $\mathcal{G}_{\text{train}}$ is the graph used to train the model, $\mathcal{E}_{\text{dev}}$ is the ground truth positive edges in the dev set inputs. The formulation above means that the model maximizes the probability of accurately predicting edges on the dev set, while observing certain training graph $\mathcal{G}_{\text{train}}$.
3.2. Recommendation with Data Augmentation
We study a simple and effective approach to augment the graph, which adds new nodes and edges to augment the original graph $\mathcal{G}$ into a larger graph $\mathcal{G}\cup\mathcal{G^{+}}$. The training objective then becomes:
(2) $\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}}\cup\mathcal{G}^{+})=\mathcal{E}_{\text{dev}})$
The augmented graph $\mathcal{G}^{+}$ is comprised of a unified set of intent nodes and heterogeneous connections to existing graph nodes, which is visualized in Figure 2 and to be described in Section 4.1.
Figure 2. IKGR’s graph augmentation steps to build intent knowledge graph with LLM. IKGR overcomes knowledge gap via in-context learning, avoids synthetic noise by focusing on the simple node-level intent entity retrieval task, and being light weight and tuning-free.
3.3. Knowledge Graph Convolution Layer
A Knowledge Graph (KG) is a directed graph composed of subject-property-object triple facts. Each triplet $(e_{h},e_{t},r)$ denotes a relationship $r$ from head entity $e_{h}$ to tail entity $e_{t}$.
Similar to KGCN (Wang et al., 2019), we employ a knowledge graph convolution layer to capture structural proximity among entities in a knowledge graph. The model learns node embeddings $\bm{E}\in R^{T_{R}\times d}$ and relation embeddings $\bm{R}\in R^{N\times d}$, where $T_{R}$ is number of relation types, $N$ is number of nodes, and $d$ is the embedding dimensionality. Denote the input features for some node $v$ as $\mathbf{v}$, the set of entities directly connected to $v$ as $\mathcal{S}_{v}$, then the output embedding is the summation-aggregated neighborhood entity and relation embeddings:
(3) $\mathbf{v}^{\text{out}}=\sigma\left({\bf W}\cdot[{\bf v}+\text{softmax}(\bm{R}[\mathcal{S}_{v}]\bm{E}[\mathcal{S}_{v}]^{T})\bm{E}[\mathcal{S}_{v}]]+{\bf b}\right)$
where $\bm{R}[\mathcal{S}_{v}]$ and $\bm{E}[\mathcal{S}_{v}]$ are the relationship embedding vector set and entity embedding vector set, each containing $|\mathcal{S}_{v}|$ of $d$-dimensional vectors. $\bf W$ and $\bf b$ are transformation weights and a bias term, respectively, and $\sigma$ is the activation function.
3.4. Knowledge Graph Embedding Translation
Due to the incomplete nature of KGs, KG completion is often leveraged as a self-supervised learning task, which predicts the missing entity $e_{h}$ or $e_{t}$ for a triplet $(e_{h},e_{t},r)$. To this end, TransE (Bordes et al., 2013), a popular knowledge graph embedding model commonly used for KG completion, enforces a translation in the embedding space: $\mathbf{e}_{h}+\mathbf{e}_{r}\approx\mathbf{e}_{t}$. This embedding translation is achieved via the following training objective: $\min\mathop{\Sigma}\limits_{\mathbf{e}_{h},\mathbf{r},\mathbf{e}_{t}}\parallel\mathbf{e}_{h}+\mathbf{r}-\mathbf{e}_{t}\parallel$.
One issue with TransE embedding is that a single relation type may correspond to multiple head entities or tail entities, leading to significant 1-to-N, N-to-1, and N-to-N issues (Wang et al., 2014). As an improvement, TransH (Wang et al., 2014) learns different representations for an entity conditioned on different relations. It assumes that each relation owns a hyperplane, and the translation between head entity and tail entity is valid only if they are projected to the same hyperplane. It defines an energy score function for a triplet as follows:
(4) $f(\mathbf{e}_{h},\mathbf{e}_{t},\mathbf{r})=\parallel\mathbf{e}_{h}^{\bot}+\mathbf{r}-\mathbf{e}_{t}^{\bot}\parallel$
where a lower score of $f(e_{h},e_{t},r)$ indicates that the triplet is likely valid. $\mathbf{e}_{h}^{\bot}$ and $\mathbf{e}_{t}^{\bot}$ are projected entity vectors:
(5) $\mathbf{e}_{h}^{\bot}=\mathbf{e}_{h}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{h}\mathbf{w}_{r}$
(6) $\mathbf{e}_{t}^{\bot}=\mathbf{e}_{t}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{t}\mathbf{w}_{r}$
where $\mathbf{w}_{r}$ and $\mathbf{r}$ are two learned vectors for a specified relation $r$. $\mathbf{w}_{r}$ denotes the projection vector of its corresponding hyperplane where $\mathbf{r}$ is the relationship embedding vector.
Finally, the training of TransH encourages the discrimination between valid triplets and incorrect ones using margin-based ranking loss. Similar to KTUP (Cao et al., 2019), we employ:
(7) $\mathcal{L}_{KG}=-\sum_{\begin{subarray}{c}(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})\in\mathcal{KG}\\ (\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})\in\mathcal{KG}^{-}\end{subarray}}\log\sigma[f(\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})-f(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})]$
where $\mathcal{KG}^{-}$ contains corrupted triplets constructed by randomly sampling a tail entity and relation. In practice, weight decay and normalization-enforcing losses are also applied to prevent overfitting.
4. Methodology
In this section, we address the following key questions regarding LLM-based knowledge graph augmentation: Q1: How to enable node-level intent entity extraction with RAG? Q2: How to inject the learned intents as structural priors into KG modeling? Q3: How to improve scoring with intent awareness at test time?
4.1. Intents extraction with RAG (Q1)
Despite significant progress in incorporating side information into recommendation systems, introducing low-quality side information may even undermine what little signal we can glean from sparse interactions. To address this challenge, the proposed IKGR focuses on an effective and simple node-level graph augmentation. We leverage a LLM to build a knowledge graph with an additional type of node: the interaction intent entities. The intent nodes are linked to the existing graph via two types of edges: exact intent $\mathcal{E}_{E}$ and related intent $\mathcal{E}_{R}$.
In order to address challenges posed by sparse user behavior scenarios, we leverage the LLM to perform data augmentation. Unlike existing LLM recommendation methods that directly synthesize user-item interactions (Wei et al., 2024), we take a more conservative approach, tasking the LLM with a simpler, more reliable role to minimize noise in the generated outputs. This approach leverages the LLM’s pre-trained common sense knowledge for accurate distillation.
We note that a forward call to the LLM with appropriate instructions is able to extract an intent entity set. For example, for item node $i$, we denote the set of intent entities as $\Omega_{i}$, $\Omega_{i}=LLM($###Item Description: {Item i description} Return a list of entities mentioned in the Item Description that the user may have intents to interact with)))). However, there are two gaps of this naive application: (1) the existence of domain-specific knowledge may make it difficult for the LLM to fully understand the context; (2) the connectivity of the extracted intents could follow a very long-tail distribution, meaning that the majority of intent nodes might only link to few user/item nodes. To mitigate these challenges, we propose the following improvements over the naive forward call.
4.2. In-context Learning
To keep the resource budget lightweight, we curate an in-context learning process without fine-tuning. For the enterprise use-case, which differs from the public dataset use-case, user profiles and item descriptions often contain abbreviations and domain-specific concepts unknown to LLMs. To address this knowledge gap, we curate a knowledge base of key-value pairs, where keys are the domain-specific abbreviations/concepts, and values are their explanations. We identify any such terminology present in the user profile or item description and append their corresponding explanations to the prompt under a ###$\#\#\#Concept\ Explanation$ tag. This enriches the context for the LLM, enabling better understanding without the need for fine-tuning.
4.3. Intent connectivity enrichment via RAG
A two-round process extracts intents and builds a knowledge graph. In the first round, a simple prompt template is used to generate specific entities, linking users and items to intent nodes. However, this initial graph may be sparsely connected, with many intent nodes linked to only a few users/items. To address this, a second round enriches connectivity by linking additional user/item nodes to existing intent nodes, avoiding the computationally expensive alternative of grouping similar intents ($\mathcal{O}(N^{2})$ complexity)
In the second round, each user and item node is connected to additional intent entities from the fixed pool of intents generated in the first round. We call these new connections ”related intents”. During this second round of related intent selection, the existing intents for a given user/item node are excluded from the retrieval step.
Denote the intent entities obtained after the first round extraction as $\bar{\Omega}$. Given a user profile or an item description $text$, the second round intent extraction prompt can be formulated as: $LLM($###Knowledge Context: … ### Options: $R(x,\mathcal{N}\backslash\Omega_{x},K)$ What are the intents mentioned in $x$ that are the most relevant?)))) , where $R(x,\bar{\Omega}\backslash\Omega_{x}^{\prime},K)$ retrieves a group of $K$ intents that are semantically similar to the input text $x$, yet have not been extracted during the first round output ($\Omega_{x}^{\prime}$).
New users and items undergo this two-round extraction to connect them to the existing, well-connected knowledge graph. Intent construction is always applied to items (representing an item satisfying a user intent). When available, user profile data is also used for intent construction, and the resulting user and item intent nodes are merged using case-insensitive exact matches. Both rounds prompt the LLM for structured output.
Figure 3. The graph module of IKGR. After constructing the knowledge graph, the embeddings of user, item and intents are fed into the knowledge graph convolution layer followed by an embedding translation layer.
4.4. Intent Translation GNN (Q2)
Having obtained the intent nodes from items and users, we train a knowledge graph-based GNN to predict user-item interaction probabilities. The architecture is shown in Figure 3, which is a knowledge graph convolution layer followed by a transH layer. We note that this component of the system can be swapped for any KG-compatible graph learning method. We posit that as long as the node-level features are efficiently decoupled and connections are made, semantic meaning can be effectively captured even with a relatively simple graph model. Therefore, we leverage a simple embedding translation-based approach to verify that building intent-item and intent-user connections helps enable more accurate user-item interaction predictions even with simpler architectures. This GNN module is implemented for the ease of the benchmarking and ablation study, so that the gain from the intent graph augmentation can be readily observed. Next, we discuss the embedding translation mechanisms and how we connect the user-item preference translation with the newly added intent embeddings at inference time.
First, we leverage a pre-trained natural language encoder to process textual node features into embeddings as input to IKGR. Given the raw embedding input, we apply a KG convolution layer described in Equation 3 to inject the intent signal into user and item nodes to prepare for interaction prediction. Then, the KG embedding translation layer is applied to encourage semantic meaning alignment given the identified intent nodes in the graph.
In the intent-augmented knowledge graph, there are three main types of relations: user possesses intent, item satisfies intent, and user consumes item. For the first two relation types, we leverage Equation 7 to train relation representation $\mathbf{r}$ and projection vector $\mathbf{w}_{r}$ without modification. For the third type, user consumes item, the projection vector $\mathbf{w}_{r}$ is still independently trained, but the translation vector $\mathbf{r}$ is based on intent node embeddings.
The motivation behind leveraging the intent nodes to compute relation representation is that both the user and the item are decoupled into lists of intents, and when there are shared or similar intent nodes, their difference is expected to be small, i.e., $\mathbf{e}_{h}^{\bot}\approx\mathbf{e}_{t}^{\bot}$ or $\parallel\mathbf{r}\parallel\approx 0$. In this way, the intent embedding introduces direct insight for the relation embedding translations. We hence build the relation embeddings $\mathbf{r}^{u,i}\in R^{d}$ as follows.
Denote the intent embeddings for user node $u$ and item node $i$ as $\mathbf{Z}^{u}\in R^{|\mathcal{S}(u)|\times d}$ and $\mathbf{Z}^{i}\in R^{|\mathcal{S}(i)|\times d}$, where $\mathcal{S}(u)$ and $\mathcal{S}(i)$ are the intent node neighbor sets.
We first calculate two matrices: $\mathbf{P}^{u,i}\in R^{|\mathcal{S}(u)||\mathcal{S}(i)|\times 1}$, containing cosine similarities between each pair of rows in $\mathbf{Z}^{u}$ and $\mathbf{Z}^{i}$, and $\mathbf{D}^{u,i}\in R^{|\mathcal{S}(u)||\mathcal{S}(i)|\times d}$, containing $\mathbf{Z}^{i}_{[q,:]}-\mathbf{Z}^{u}_{[p,:]}$ for each row index $p=0,1,2,\cdots|\mathcal{S}(i)|-1$ and $q=0,1,2,\cdots|\mathcal{S}(u)|-1$.
Then, to enforce the relationship of more similar intent pairs inducing a smaller translation vector, the resulting vector $\mathbf{r}^{u,i}$ is computed via:
(8) $\mathbf{r}^{u,i}=\text{softmax}(\mathbf{P}^{u,i})^{T}\mathbf{D}^{u,i}$
4.5. Intent Aware Scoring (Q3)
To further leverage signal from the intents extracted during KG construction, we incorporate them into the user-item interaction scoring function, in conjunction with the more traditional embedding similarity score. At inference time, given a tuple of user and item $(u,i)$, we derive embedding matching and intent matching scores between a user and an item, and the final score is a hybrid combination of them. The embedding matching score between $u$ and $i$ is the cosine similarity between their embeddings:
(9) $z_{u,i}=\frac{\mathbf{e}_{u}\cdot\mathbf{e}_{i}}{\parallel\mathbf{e}_{u}\parallel\cdot\parallel\mathbf{e}_{i}\parallel}$
We denote a single intent extracted from $q$ as $n^{(q)}$, and the collection of all intents for $q$ as $\Omega_{q}$. Similarly, a single intent extracted from $u$ and their collections is denoted as $n^{(u)}$ and $\Omega_{u}$, and a single intent extracted from $i$ and their collections is denoted as $n^{(i)}$, and $\Omega_{i}$. Additionally, $\mathbf{e}_{x}$ is used to represent the embedding of some entity $x$.
The intent matching score between $u$ and $i$ is determined by:
(10) $y_{u,i}=\mathop{max}\limits_{n^{(u)}\in\Omega_{u},n^{(i)}\in\Omega_{i}}\frac{\mathbf{e}_{n^{(u)}}\cdot\mathbf{e}_{n^{(i)}}}{||\mathbf{e}_{n^{(u)}}\parallel\cdot\parallel\mathbf{e}_{n^{(i)}}\parallel}\cdot 0.5^{\mathbb{I}(\Omega_{u}\cap\Omega_{i}=\varnothing)}$
Finally, the score for the triplet $(u,i)$ is the hybrid mixture of the embedding score and intent matching score:
(11) $score(u,i)=y_{u,i}+\lambda z_{u,i}$
where $\lambda$ is an empirical mixer coefficient.
4.6. Inference Procedure And Complexity Analysis
5. Experiments
In this section, we verify the performance of the proposed recommender using both proprietary and open-source data. For the proprietary data, we introduce IKGR to an enterprise knowledge search platform to test IKGR’s effectiveness in re-ranking search results. We also benchmark against existing baselines on four open-source recommendation datasets to verify the results in diverse scenarios. The details of these datasets are presented in Table 1.
We use Llama-3.1-8B model for LLM inference, and all-mpnet-base-v2 from sentence transformer for encoding textual features. We retrieve the top 100 prebuilt intent candidates in the kNN retrieval step of RAG. We apply an 8:1:1 ratio when sampling the positive/negative edges for train/dev/test sets in the graph. The coefficient $\lambda$ in Equation 11 is set to 0.1 based on grid search result on the validation set.
[…]
Datasets Search Beauty Books Steam Yelp2022
#Users 36 0333603336\,03336 033 40 2264022640\,22640 226 251 394251394251\,394251 394 281 428281428281\,428281 428 1 987 89819878981\,987\,8981 987 898
#Items 872 678872678872\,678872 678 54 5425454254\,54254 542 25 6062560625\,60625 606 13 0441304413\,04413 044 150 347150347150\,347150 347
#Inter 3.5M 0.35M 3.2M 3.5M 6.9M
Density 0.011% 0.02% 0.05% 0.095% 0.002%
#Intents 495 285495285495\,285495 285 39 3053930539\,30539 305 45 9324593245\,93245 932 53 9405394053\,94053 940 65 0406504065\,04065 040
#IntEdges 6.9M 209K 390K 231K 409K
AvgIntDeg 13.9 5.3 8.5 4.3 6.3
5.1. Baselines and Datasets
Enterprise Search. In the enterprise search setting, a set of developers query the search engine to search for datasets published by other developers within the enterprise. The search engine first uses BM25 (Whissell and Clarke, 2011) to retrieve a list of item candidates, then the IKGR is applied to rerank the search results. Evaluations are done over user’s feedback on how high the item of interest could be ranked among the final reranked list. In this setting, the items are enterprise datasets, which consist of text features such as dataset name, description, column names, ID labels, etc. The users are all registered within the enterprise and their developer profiles could be collected through separate channels to offer a hint about their dataset consumption preference. The IKGR is trained over historical user-dataset consumption interactions collected from separate channels. This dataset is labeled as Search in dataset description table and result tables.
Books, Beauty. These datasets are obtained from Amazon review [^1] http://jmcauley.ucsd.edu/data/amazon/ in (McAuley et al., 2015), which contains a variety of categories. We utilize the Books and Beauty categories. We leverage the features of title, sales type, sales rank, categories, price, and brand.
Steam [^2] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data. This is a dataset collected from Steam (Kang and McAuley, 2018), a large online video game distribution platform. We leverage the item features of app name, genres, publisher, sentiment, specs, tags.
Yelp2022333https://www.yelp.com/dataset. This is a popular dataset for business recommendation. Given the large size, we use the transaction records after January 31st, 2022. We treat the categories of businesses as attributes for items, and user compliment types as attributes for users.
To evaluate the ranking list of all the models, we employ Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR). Considering we only have one ground truth item for each user, HR@$k$ is equivalent to Recall@$k$ and proportional to Precision@$k$; MRR is equivalent to Mean Average Precision (MAP). We report HR and NDCG with $k={1,5,10}$. For all these metrics, the higher the value, the better the performance.
To verify the effectiveness of our method, we compare it with the following representative baselines:
POP: It is the simplest baseline that ranks items according to their popularity judged by the number of interactions.
BPR-MF (Rendle et al., 2012): It optimizes the matrix factorization with implicit feedback using a pairwise ranking loss.
FM (Rendle, 2010): The factorization machine models all interactions between variables using factorized parameters.
ConvNCF (He et al., 2018): It employs a convolutional neural network to learn high-order correlations among embedding dimensions neural collaborative filtering.
FPMC (Rendle et al., 2010): It captures users’ general taste as well as their sequential behaviors by combing MF with first-order Markov chains.
RippleNet (Wang et al., 2018): It uses a single layer of embedding translation by applying transH to user-item interaction.
KGCN (Wang et al., 2019): The KG based graph convolutional networks captures inter-item relatedness effectively by mining their associated attributes on the KG.
KTUP (Cao et al., 2019): It uses a single layer of embedding translation by applying TransH to user-item interaction.
Table 2. Performance comparison of different methods. Bold scores are the best in each row, while underlined scores are the second best.
Datasets Metric POP BPR-MF ConvNCF FPMC RippleNet KGCN KTUP IKGR
Search HR@1 0.0001 0.0041 0.0075 0.0048 0.0024 0.0051 0.0078 0.0086
HR@5 0.0023 0.0100 0.0149 0.0125 0.0118 0.0161 0.0166 0.0202
HR@10 0.0076 0.0174 0.0262 0.0225 0.0218 0.0232 0.0262 0.0267
NDCG@5 0.0029 0.0068 0.0090 0.0074 0.0068 0.0101 0.0137 0.0151
NDCG@10 0.0046 0.0094 0.0112 0.0107 0.0087 0.0123 0.0142 0.0172
MRR 0.0046 0.0108 0.0070 0.0097 0.0091 0.0100 0.0130 0.0153
Beauty HR@1 0.0405 0.0552 0.1005 0.1009 0.0532 0.0984 0.1783 0.1369
HR@5 0.0377 0.1844 0.2754 0.2379 0.1972 0.3120 0.3388 0.3316
HR@10 0.1074 0.3540 0.4387 0.3957 0.3695 0.4204 0.4610 0.4846
NDCG@5 0.0763 0.1245 0.1758 0.1335 0.1307 0.1729 0.2583 0.2806
NDCG@10 0.0443 0.1731 0.1961 0.2123 0.1713 0.2345 0.2814 0.2939
MRR 0.0909 0.1792 0.1736 0.1752 0.1382 0.2254 0.2581 0.2641
Books HR@1 0.0015 0.0324 0.1124 0.0620 0.0480 0.0853 0.1125 0.1251
HR@5 0.0524 0.1663 0.2248 0.1859 0.1691 0.2465 0.2652 0.3197
HR@10 0.1073 0.2649 0.3181 0.3217 0.3553 0.3582 0.3614 0.4248
NDCG@5 0.0314 0.0983 0.1371 0.1268 0.1031 0.1579 0.1921 0.2097
NDCG@10 0.0532 0.1603 0.1633 0.1703 0.1607 0.2016 0.2415 0.2814
MRR 0.0788 0.1317 0.1567 0.1532 0.1407 0.1832 0.2193 0.2672
Steam HR@1 0.0091 0.0547 0.0860 0.0526 0.0304 0.0641 0.1060 0.1095
HR@5 0.0370 0.1196 0.1976 0.1698 0.1429 0.1966 0.2445 0.2759
HR@10 0.0901 0.2379 0.3374 0.2717 0.2735 0.3218 0.3126 0.3574
NDCG@5 0.0291 0.1035 0.1121 0.1094 0.0915 0.1496 0.1614 0.1735
NDCG@10 0.0459 0.1244 0.1471 0.1384 0.1161 0.1905 0.1809 0.2212
MRR 0.0507 0.1265 0.1382 0.1272 0.1294 0.1528 0.1957 0.2168
Yelp2022 HR@1 0.0171 0.0636 0.1177 0.0524 0.0495 0.0917 0.1005 0.0989
HR@5 0.0366 0.1555 0.1987 0.1911 0.1513 0.2405 0.2395 0.2869
HR@10 0.1018 0.2462 0.3576 0.3028 0.2996 0.3395 0.3317 0.3966
NDCG@5 0.0233 0.0823 0.1395 0.1131 0.1201 0.1397 0.2067 0.2078
NDCG@10 0.0334 0.1364 0.1434 0.1527 0.1225 0.1861 0.2145 0.2277
MRR 0.0593 0.1327 0.1542 0.1200 0.1356 0.1578 0.1906 0.2100
Observing the results in Table 2, both datasets and models show statistical performance differences. From the dataset perspective, different datasets have different levels of knowledge gaps, with the Search dataset possibly having the largest knowledge gap for the LLM. This could be indicated by the number of intents and average intent node degrees in Table 1, where the Search dataset’s intent count is about one magnitude larger than other datasets. On the other hand, this is also observed by the performance difference of the IKGR model, where the same architecture is applied on different LLM augmented datasets. As IKGR performances are lower on the Search dataset, it hints that interaction prediction under the knowledge gap is more difficult.
From the approaches perspective, IKGR shows new state-of-the-art performances especially on the Search dataset with the largest knowledge gap. On the other hand, across these datasets, the knowledge graph embedding translation approaches (IKGR, TKUP) or GNN based approaches (KGCN, RippleNet) performs better than factorization machines or collaborative filtering based approaches (BPR-MF, FPMC, ConvNCF), demonstrating the gain from structural embedding learning.
5.2. Ablation Studies
Table 3. Ablation Study
Metrics HR@1 HR@5 HR@10 NDCG@5 NDCG@10 MRR
IKGR 0.0086 0.0202 0.0267 0.0151 0.0172 0.0153
Exact Int. Only 0.0077 0.0185 0.0254 0.0147 0.0163 0.0134
No Aug. 0.0073 0.0175 0.0240 0.0131 0.0150 0.0125
Inter. Aug. 0.0070 0.0166 0.0220 0.0126 0.0141 0.0117
No Int. scoring 0.0082 0.0193 0.0256 0.0140 0.0164 0.0139
No Int. Trans. 0.0078 0.0184 0.048 0.0132 0.0152 0.0130
To investigate the influence of each component over the enterprise search data, we conduct a list of experiments that drop each component in the proposed pipeline and compare with the original pipeline on the Search dataset. Results are presented in Table 3.
In Table 3, IKGR means the original approach without component dropping, Exact Int. Only means dropping the second round of related intent retrieval using RAG, and only exact intent edges are presented in the graph, without the related intent edges. No Aug. means training the architecture of IKGR, which is a naive knowledge graph convolution followed by an embedding translation layer, without adding the intent node in the graph. Inter. Aug. means to employ LLM only as a user-item implicit feedback augmentor without adding the intent nodes (Wei et al., 2024). In this setting the LLM is asked to select items that the user might be likely or unlikely to interact with. No Int. scoring means using a simple cosine similarity based score and we do not apply intent-aware scoring function as in Equation 11. No Int. Trans. means intent translation is not leveraged to compute the translation as in Equation 8.
As observed in the results shown in Table 3, the relative contributions of each components of IKGR can be estimated. The most significant observation is that the intent edges augmentation steps as a whole helps to densify the graph almost twice, as reflected by the #IntEdges and #Inter in Table 1, and the MRR performance correspondingly improved 22% (0.0125 to 0.0153) shown in Table 3. This confirms that adding intent relations to the graph improves the IKGR performance while densifying the graph with meaningful knowledge connectivities.
The proposed intent node embedding improves the performance by offering a straightforward embedding structure between user and item vectors, hence guiding the learning procedure, as shown in Table 3 that learning the relation vectors from scratch and removing the intent based scoring both harms the performance.
6. Conclusions
In this work, we propose IKGR, a knowledge graph based recommender built with a Large Language Model (LLM). The proposed method features a data augmentation step to explicitly extract entities that the users have intents to interact with, and learns node embeddings over the knowledge graph using an embedding translation layer to combine the intent structure knowledge. This work takes the enterprise search personalization as a case study, and verifies that (1) when knowledge gap exists, using a simplified node-level augmentation task helps learn embeddings, while synthesizing interactions harms the model performance and introduces noise; (2) injecting intent structure prior into the modeling helps better capturing the semantic structure and boosts embedding learning.

### v3

@@
-Explain What You Mean: Intent Augmented Knowledge Graph Recommender Built With An LLM
-Abstract.
-Interaction sparsity is a long-standing challenge in recommendation systems.
-Sparsity manifests in environments with disproportional cardinality of groupings of entities, such as users and products in an online marketplace.
-It is also found for newly introduced entities, described as the cold-start problem.
-Recent efforts to mitigate this issue either enrich the connectivity data by incorporating social networks or external knowledge graphs, or fine-tune LLMs into interaction augmenters or next-item recommenders.
-However, these techniques tend to be resource demanding, requiring high computational power.
-They also have several limitations, including data availability, low quality, or synthetic noise issues.
-In this work, we propose LLM-based Intent Knowledge Graph Recommender (IKGR), a novel framework that leverages retrieval-augmented generation and an encoding approach to construct and densify a knowledge graph.
-IKGR leverages latent user-item affinities from an interaction knowledge graph and further densifies it through mutual intent connectivity.
-This addresses sparsity issues and allows the model to make intent-grounded recommendations with an interpretable embedding translation layer.
-Through extensive experiments on real-world datasets, we demonstrate that IKGR overcomes knowledge gaps and achieves substantial gains over state-of-the-art baselines on both publicly available and our internal recommendation datasets.
-Recommendation system, Large Language Model (LLM), Knowledge Graph, Retrieval Augmented Generation (RAG)
-1. Introduction
-Recommender systems fulfill a crucial role by helping users navigate through an ever-increasing catalog of online content and services to find what is most relevant for their personalized interests (Meng et al., 2023a; Wei et al., 2022).
-In enterprise applications, recommender algorithms are also widely deployed in internal search platforms or forums to help employees and developers ease the understanding of enterprise knowledge and the discovery of helpful articles, tips, or properties (Verma et al., 2015; Roy and Dutta, 2022).
-A wide family of recommendation strategies has demonstrated success in modeling users’ historical interactions, such as collaborative filtering (CF) (Koren et al., 2021), graph-based methods (Pei et al., 2019; Liu et al., 2020; Zuo et al., 2024), and more recently recommenders based on Large Language Models (LLMs) (Kim et al., 2024; Bao et al., 2023; Wu et al., 2024).
-As these methods rely on implicit feedback from previous interactions, their performance tends to degrade for cold-starting new users or items with sparse interactions, hence it has always been a central topic in recommender systems to efficiently leverage user/item features and interactions to overcome sparse implicit feedback.
-To address the cold start problem, researchers have proposed various approaches to augment the plain user-item interaction data with additional information.
-Depending on the source of added contextual information, these augmentation methods can be grouped into three categories: (1) enhancing user/item attributes with existing data, such as item attributes (Ying et al., 2023), user-generated content (Fan et al., 2019; Meng et al., 2023b), or multi-modal features (Yi et al., 2022); (2) enhancing linkages with existing data, such as social networks (Jamali and Ester, 2010) or external knowledge graphs (Wang et al., 2018), and (3) enhancing user/item links and attributes with LLMs, such as interaction augmentation (Wei et al., 2024).
-These approaches attempt to enrich entity attributes, increase interactions, or add contextual information to help the recommender learn better representations and provide improved recommendations.
-However, these augmentation methods can still face several challenges, which we categorize as follows.
-i) Synthetic data noise caused by preference differences and knowledge gaps.
-Handling noisy data is a long-standing challenge faced by recommender systems trained on implicit feedback, where temporary or individualized user preferences can induce false positive or false negative samples.
-The noise issue can be further exacerbated by certain attempts to alleviate sparsity, such as the generation of synthetic user interactions with LLMs (Wei et al., 2024).
-LLM-generated preferences can introduce additional distribution skewness, potentially over-representing popular items, and failing to capture the long-tail distribution of real-world user preferences.
-To mitigate the additional data noise, careful calibration techniques need to be employed, e.g., loss calibration and training sample selection (Wei et al., 2024).
-Furthermore, LLMs may exhibit knowledge gaps, lacking necessary domain-specific context.
-In enterprise use cases especially, there is often a gap between private enterprise data and public corpora, leading to difficulty for both LLMs and traditional data augmentation approaches.
-ii) Excessive resource budget.
-Another line of research has leveraged LLMs as an end-to-end function to perform the recommendation task (Wu et al., 2024; Deldjoo et al., 2024; Bao et al., 2023; Kim et al., 2024).
-However, fine-tuning the LLM to mitigate the knowledge gap tends to become resource intensive both in terms of the computation budget and developer efforts to calibrate the data and the model for each specific use case.
-iii) Lack of fine-grained understanding.
-Among the LLM-based linking approaches, their predictions are mostly based on natural language representations rather than fact-based knowledge graph structures, hindering interpretability and targeted adjustments.
-For example, when recommending popular head items, it may not be clear if the recommendation is based on a genuine user affinity or simply the popularity of the item.
-A bias that is hard to correct without understanding the LLM’s underlying reasoning.
-Furthermore, as the number of real user interactions grows over time, the relative value of LLM-synthesized edges may diminish.
-Existing node enrichment methods use side information to augment node attributes with text-based features, but typically do not decompose them into graph structures, preventing the use of graph-based inference techniques (Ko et al., 2022; Liang et al., 2023).
-Figure 1.
-In enterprise search, the queries could contain special terminologies and acronyms, where traditional search engines or personalized rerankers fail to capture the real intents under such knowledge gap.
+Tuning-Free LLM Can Build A Strong Recommender Under Sparse Connectivity And Knowledge Gap Via Extracting Intent
+Abstract
+Recent advances in recommendation with large language models (LLMs) often rely on either commonsense augmentation at the item-category level or implicit intent modeling on existing knowledge graphs.
+However, such approaches struggle to capture grounded user intents and to handle sparsity and cold-start scenarios.
+In this work, we present LLM-based Intent Knowledge Graph Recommender (IKGR), a novel framework that constructs an intent-centric knowledge graph where both users and items are explicitly linked to intent nodes extracted by a tuning-free, RAG-guided LLM pipeline.
+By grounding intents in external knowledge sources and user profiles, IKGR canonically represents what a user seeks and what an item satisfies as first-class entities.
+To alleviate sparsity, we further introduce a mutual-intent connectivity densification strategy, which shortens semantic paths between users and long-tail items without requiring cross-graph fusion.
+Finally, a lightweight GNN layer is employed on top of the intent-enhanced graph to produce recommendation signals with low latency.
+Extensive experiments on public and enterprise datasets demonstrate that IKGR consistently outperforms strong baselines, particularly on cold-start and long-tail slices, while remaining efficient through a fully offline LLM pipeline.
+1 Introduction
+Modern recommender systems are expected to reason over sparse and evolving interactions while serving highly personalized needs across vast catalogs.
+This challenge is acute in enterprise environments, where recommendations support internal search and knowledge discovery yet must cope with heterogeneous vocabularies, domain jargon, and long-tail content [1, 2].
+Collaborative filtering and graph-based models have improved user–item representation learning [3, 4, 5, 6], while knowledge-aware recommenders further leverage structured relations for interpretability [7, 8].
+Nevertheless, their effectiveness is bounded by incomplete coverage and weak connectivity, particularly for long-tail and cold-start cases [9].
+To mitigate incomplete coverage and weak connectivity, a prominent direction uses LLMs to inject commonsense relations, typically complement or substitute links then fuses them with an existing item KG (CSRec) [10].
+This line is attractive because (i) it regularizes sparse graphs with priors that are broadly valid (jackets complement sweaters; lenses complement cameras), (ii) it is offline-friendly (relations are generated once and reused), and (iii) it improves item-side coverage where merchant metadata is incomplete.
+However, three limitations recur in practice.
+(1) Granularity & intent mismatch.
+Category-level commonsense smooths the space but only loosely correlates with user-specific intent.
+Two users who click the same camera page may have very different intents (e.g., low-light astrophotography vs.
+lightweight travel kit).
+Category edges cannot capture these fine distinctions, which limits cold-start personalization.
+(2) Cross-graph fusion & alignment risk.
+CSRec-style pipelines typically align an LLM-augmented commonsense graph with an existing metadata/interaction graph.
+This requires ontology matching, entity resolution, and confidence thresholds.
+Errors here introduce structural noise that is hard to debug, and different domains require repeated re-alignment [8].
+(3) Temporal & domain drift.
+Off-the-shelf commonsense tends to be domain-neutral and slow-moving.
+Enterprise or fast-evolving domains (e.g., internal tools, niche APIs) quickly deviate; commonsense edges may become stale or irrelevant without domain grounding [11, 12].
+Another line remains within a pre-existing KG and learns latent mixtures over relations to explain interactions [13, 14, 15].
+This method has several strengths: it’s entirely symbolic and structural, avoiding the need for LLMs; it uses relation paths to capture higher-order semantics; and it doesn’t require cross-graph fusion because it operates within a single KG.
+However, these pre-existing KG methods have several notable gaps.
+First, the intent remains latent and difficult to audit, as it’s represented as a vector or a mixture over relations rather than an explicit, human-readable node.
+This limits explainability and hinders downstream applications like intent analytics.
+Additionally, the model is bounded by the existing KG, meaning that if the base KG is sparse or doesn’t align with how users actually express themselves, the model struggles to capture missing semantics and often overfits to popular hubs, which hurts its performance on the long tail [9].
+Finally, there is a weak tie to external knowledge, which means issues like polysemy and synonymy in relation labels (e.g., guide, how-to, playbook) and unstandardized domain jargon persist, as the model lacks explicit text grounding or retrieval capabilities [16].
+A third family places the LLM directly in the loop — either to synthesize interactions (augment clicks/purchases) or to produce rankings end-to-end [17, 18, 19].
+This approach offers benefits such as broader semantic coverage and simplified modeling with fewer bespoke modules.
+However, the trade-offs are substantial.
+Inference-time LLM calls introduce significant latency and cost, complicating service-level agreements and A/B testing at scale, especially for multi-stage recommenders [12].
+Furthermore, synthetic interactions often lead to distribution shift and popularity bias, as they over-represent head items and reflect the LLM’s generic priors rather than the platform’s actual demand patterns; even with loss calibration, realistic long-tail fidelity is difficult to guarantee [19].
+Finally, governance and reproducibility are challenging due to frequent model or LoRA updates that change behavior and varying privacy filtering across deployments, making it difficult to maintain offline-online consistency and respond to incidents in enterprise environments [11].
+Figure 1: In enterprise search, the queries could contain special terminologies and acronyms, where traditional search engines or personalized rerankers fail to capture the real intents under such knowledge gap.
 IKGR addresses the challenge via injecting fine-grained understanding to textual features.
-Additionally, these approaches may generate abbreviated textual descriptions, which can impair the model’s representations and subsequent recommendations.
-In order to build a recommender that is good at handling the knowledge gap mentioned above, this work proposes to leverage the general reasoning ability of the LLM to augment the interactions.
-We take the the enterprise search as a primary case study as visualized in Figure 1, while keeping our approach applicable to more general scenarios.
-Our objective is to address the issue of noisy LLM generations under knowledge gap while maximally reducing the computation budget.
-We propose Intent Knowledge Graph Recommender (IKGR), an LLM-based pipeline, to explicitly augment node-level information.
-We build a knowledge graph that is naturally linked to the user-item interaction graph, as a way to enhance user/item node side information.
-Inspired by their impressive natural language understanding ability, we utilize LLMs to augment the interaction graph.
-We also simplify the task for the LLM to only boost the node-level understanding, so that more reliable results can be expected.
-In the proposed IKGR pipeline, the LLM is asked to perform a node-level retrieval task, which returns a list of entities that reflect users’ consumption intents.
-To overcome the knowledge gap, we carefully prepare an external knowledge base, explaining common abbreviations and terms.
-To reduce the computation budget, we design the framework as tuning-free, while retrieving the knowledge explanations from the knowledge base and feed to the LLM via in-context learning.
-We also design the backend method to be a lightweight Graph Neural Network (GNN) based on embedding translation, which naturally leverages the learned intent representation to inject structure prior into the graph.
-As the main contribution of this work, we see two major advantages of using an LLM to augment the knowledge graph in the IKGR pipeline.
-Firstly, IKGR embraces the shift from an heuristic embedding learning structure to an LLM-based explicit knowledge extraction regime.
-It leverages the LLM to predict a new type of node: the interaction intent from a natural language perspective.
-Unlike previous approaches that rely on uninterpretable vector embeddings to capture multifaceted node attributes, IKGR recognizes that sub-node intents written in natural language format are often extractable from node features based on LLM.
-The natural language representation provide a more intuitive reflection of user preferences and interaction intents compared to indirect embeddings, and could be more flexible in terms of intent quantities.
-By incorporating LLMs, IKGR captures the richness and context of natural language, enhancing the accuracy and effectiveness of recommendations.
-Secondly, the low-quality and incomplete side information is enhanced by leveraging the extensive knowledge of LLMs gained through the pre-training stage, which allows for a tuning-free model that is both aware of vast real-world knowledge and understands user preferences and interaction intents.
-This integration prevents the dispersion of features across separate vector spaces and provides more accurate results.
-  • We tackle the data sparsity challenge in recommendation systems, and propose IKGR, a knowledge graph augmentation approach with the LLM.
-IKGR leverages in-context learning with LLM to mitigate knowledge gap, and focuses on the node-level intent understanding and linkage to avoid introducing noise in synthesizing interactions.
-  • The proposed IKGR is LLM-tuning-free, and is able to build knowledge graph inherently linked to the user-item interactions without the need for entity alignment.
-IKGR further leverages the learned intent nodes to inject structural prior in the embedding space with a knowledge graph translation layer.
-  • Results on the enterprise search dataset demonstrate that the proposed IKGR effectively mitigate the knowledge gap, and outperforms the state-of-the-art baselines on search engine optimization reranking.
-Results on open-sourced datasets also showed the superiority of IKGR on the general recommendation scenario.
-2. Related Works
-Personalized Re-ranking Recommender System.
-Re-ranking in recommender systems aims to refine an initially ranked list to better capture user preferences and item relationships.
-Traditional ranking methods optimize global performance but often overlook the mutual influence between items and user-specific intent (Pei et al., 2019).
-Recent approaches address these limitations by integrating semantic tag information (Liang et al., 2010), multi-stage ranking optimization (Evnine et al., 2024), and item relationships (Liu et al., 2020).
-Transformer-based models effectively model global item interactions, while graph-based methods leverage item relationships for improved ranking (Pei et al., 2019; Liu et al., 2020).
-Furthermore, self-supervised learning enhances sequential recommendation by mitigating data sparsity issues (Zhou et al., 2020).
-These methods collectively demonstrate that incorporating user personalization, item dependencies, and efficient ranking strategies significantly enhances re-ranking effectiveness.
+Summarizing the pros and cons of various existing methods, we distill three core challenges to LLM and KG based recommenders.
+(i) User intent extraction is an effective way to clarify and densify the KG, but this step needs to be more grounded.
+User profiles, queries, and enterprise documents are messy: synonyms, abbreviations, internal codenames, polysemy, and multi-lingual snippets all coexist.
+Extracting what a user seeks and what an item satisfies requires more than NER; it needs disambiguation (MLP = multilayer perceptron vs.
+marketing launch plan), normalization to a stable vocabulary, and aggregation across sources.
+Pure LLM prompts help but are prone to hallucination and label drift across batches.
+Without external knowledge grounding (retrieval of glossaries, wikis, policy pages) and schema-aware canonicalization, intent nodes will be noisy, redundant, and unmaintainable [16, 11].
+This suggests the approach to elevate intent to first-class nodes extracted and normalized via RAG-guided LLMs.
+(ii) The extracted intent needs to be more directly aligned with the existing user and items.
+The goal is to shorten semantic paths by adding edges that reflect shared or similar intents (user $\leftrightarrow$ intent, item $\leftrightarrow$ intent), thereby enabling information to flow even when user–item links are missing.
+This step is better achieved by leveraging the extracted intents to densify a single graph, rather than through a separate cross-graph alignment stage.
+If densification requires merging a separate LLM commonsense graph with an interaction/metadata graph, we re-introduce entity resolution and ontology mismatch—exactly the failure modes that cause silent errors in production [10, 8].
+(iii) The system should be efficient, stable, and scalable for deployment.
+An efficient system should keep all heavy LLM work offline.
+That means handling batch extraction, incremental refresh, and backfill for new content/users; caching retrieval results; and defining compatible online components (e.g., a small GNN layer) [12].
+To avoid synthetic-data drift [19], improvements should also come from structural connectivity and explicit intents, not from generating pseudo-interactions.
+In response to these needs, we propose LLM-based Intent Knowledge Graph Recommender (IKGR), an intent-centric knowledge graph framework.
+First, we instantiate intent nodes and link users and items to them via a tuning-free, RAG-guided LLM extraction pipeline, grounded in external knowledge [16, 11].
+Second, we introduce mutual-intent connectivity densification, which structurally improves graph connectivity for sparse and cold-start regimes without relying on cross-graph fusion [10, 18].
+Finally, a lightweight GNN layer learns over the intent-enhanced graph to produce low-latency recommendations, fully decoupled from online LLM inference [17, 12].
+Empirically, IKGR consistently outperforms strong baselines on public and enterprise datasets, especially on long-tail slices, while providing interpretable pathways via explicit intents.
+The contributions of IKGR can be summarized as follows.
+  • IKGR introduces an intent-centric KG construction approach that turns user and item intents into first-class nodes.
+Using a tuning-free, RAG-guided LLM extractor, it canonically links users and items to intents with high precision.
+  • IKGR involves an intent connectivity densification step, which shortens semantic paths between users and long-tail items, improving sparsity and cold-start performance without cross-graph fusion.
+  • IKGR’s pipeline is offline and low-latency online by construction, and consistently outperforms state-of-the-art KG or LLM baselines on public and enterprise datasets, with notable improvements on long-tail and cold-start slices.
+2 Related Works
+LLM and KG based Recommenders.
+Recent responses to sparsity and cold start fall into four camps.
+(1) LLM-assisted item-side augmentation injects category/attribute-level commonsense (e.g., complement/substitute) and fuses it into an item KG (CSRec) [10].
+This regularizes sparse graphs with broadly valid priors and is offline-friendly, yet it struggles to align with user-specific intent (two camera clicks can imply very different needs), requires brittle cross-graph ontology/entity alignment, and drifts in fast-moving or enterprise domains without domain grounding [8, 11, 12].
+(2) KG-based implicit-intent modeling (KGIN) stays inside a single KG and explains interactions via latent mixtures over relations [13].
+It avoids fusion and can exploit multi-hop paths, but the intent remains a non-auditable latent vector, performance is bounded by the coverage of the base KG (hurting the long tail), and weak ties to text/external knowledge leave synonymy/polysemy unresolved [9, 16].
+(3) LLM-as-Recommender / synthetic interactions broaden semantic coverage or even replace ranking end-to-end [17, 18, 19], but introduce inference-time cost and tail latency, amplify popularity bias in generated data, and complicate governance and reproducibility under frequent model updates [12, 11].
+(4) Text/multimodal node enrichment improves representations by attaching reviews/UGC/features to nodes, yet leaves knowledge ungraphized—no new edges/nodes for message passing or structural densification [20, 21].
+Taken together, current trends either inject coarse item-side priors, keep intent latent and hard to use beyond ranking, or pay online-LLM costs and stability penalties.
 Recommender Approaches Discovering Intent.
 Recent works explore intent modeling by leveraging external signals such as search queries and knowledge graphs.
-UDITSR (Zhang et al., 2024) jointly models search and recommendation by utilizing explicit search queries to infer implicit demand intents, while a dual-intent translation mechanism captures relationships between inherent intent, demand intent, and item interactions.
-Alternatively, knowledge graph-based methods, such as KGIN (Wang et al., 2021), refine intent discovery by representing user-item interactions through fine-grained relational paths, improving interpretability and recommendation quality.
+UDITSR [22] jointly models search and recommendation by utilizing explicit search queries to infer implicit demand intents, while a dual-intent translation mechanism captures relationships between inherent intent, demand intent, and item interactions.
+Alternatively, knowledge graph-based methods, such as KGIN [23], refine intent discovery by representing user-item interactions through fine-grained relational paths, improving interpretability and recommendation quality.
 These approaches demonstrate that integrating external intent signals and structured relational modeling enhances intent-aware recommendations.
-Recommender Approaches Addressing Data Sparsity.
-Recent works tackle the data sparsity issue through enhanced model architectures, graph-based techniques, and knowledge-aware methods.
-The Item History Model (IHM) improves cold-start item recommendations by directly injecting user-interaction data into the item tower and employing an inductive structure for real-time inference (Zuo et al., 2024).
-Cold Brew addresses sparsity in graph neural networks by distilling node representations, mitigating the impact of missing or noisy neighbors (Zheng et al., 2021).
-RippleNet further alleviates sparsity by propagating user preferences through knowledge graph relations, enriching item representations beyond collaborative signals (Wang et al., 2018).
-These approaches demonstrate that leveraging historical interactions, distilling structural knowledge, and integrating external information effectively mitigate data sparsity in recommendation systems.
 LLM-Based Interaction Augmentation for Recommender Systems.
-Gen-RecSys provides a comprehensive overview of generative models in recommendation, highlighting LLM-driven natural language understanding and multimodal integration (Deldjoo et al., 2024).
-LLMRec introduces graph augmentation strategies using LLMs to enrich interaction graphs, refine side information, and denoise implicit feedback, demonstrating performance gains across benchmark datasets (Wei et al., 2024).
-BLAIR further bridges language and recommendation by pretraining sentence embeddings on large-scale review data, improving item retrieval in complex natural language contexts (Hou et al., 2024).
-Additionally, BERT4Rec employs bidirectional self-attention to model user behavior sequences, overcoming limitations of traditional sequential models (Sun et al., 2019).
+Gen-RecSys provides a comprehensive overview of generative models in recommendation, highlighting LLM-driven natural language understanding and multimodal integration [24].
+LLMRec introduces graph augmentation strategies using LLMs to enrich interaction graphs, refine side information, and denoise implicit feedback, demonstrating performance gains across benchmark datasets [19].
+BLAIR further bridges language and recommendation by pretraining sentence embeddings on large-scale review data, improving item retrieval in complex natural language contexts [25].
+Additionally, BERT4Rec employs bidirectional self-attention to model user behavior sequences, overcoming limitations of traditional sequential models [26].
 These studies collectively showcase the potential of LLMs in augmenting interactions, refining representations, and enhancing recommendation performance.
-Directly Leveraging LLM as the Recommender.
-A-LLMRec enhances LLM-based recommendation by leveraging embeddings from state-of-the-art collaborative filtering models, excelling in both cold and warm scenarios while maintaining efficiency and model-agnostic integration (Kim et al., 2024).
-CoRAL introduces collaborative retrieval-augmented prompting, addressing LLMs’ reliance on semantic information by incorporating user-item interactions through reinforcement learning-based retrieval policies, significantly improving long-tail recommendation (Wu et al., 2024).
-TALLRec proposes a tuning framework to align LLMs with recommendation-specific tasks, demonstrating strong generalization and efficiency even with limited training data (Bao et al., 2023).
-These approaches collectively highlight the potential of LLMs as standalone recommenders by addressing cold-start challenges, enhancing collaborative reasoning, and improving alignment with recommendation-specific objectives.
-3. Preliminaries
-This work brings techniques from recommendation systems, knowledge graph-based GNNs, and Retrieval Augmented Generation (RAG) with LLMs.
-To help readers better digest our terminology and approaches, we leverage this section to present the problem formulation and commonly adopted techniques across these domains, and introduce our contributions in Section 4.
-3.1. Graph Formulation Of Recommendation System
+More related works regarding data sparsity, LLM as recommender, and personalized re-rankers are detailed in Appendix C.
+3 Preliminaries
+We leverage this section to present the problem formulation and commonly adopted techniques across these domains, and introduce our contributions in Section 4.
+Figure 2: IKGR’s graph augmentation steps to build intent knowledge graph with LLM.
+IKGR overcomes knowledge gap via in-context learning, avoids synthetic noise by focusing on the simple node-level intent entity retrieval task, and being light weight and tuning-free.
+3.1 Graph Formulation of Recommendation System
 We consider the recommendation system where a set of registered users $\mathcal{U}$ interact over a set of items $\mathcal{I}$.
@@
 In mathematical formulation, the objective of the recommendation systems is to maximize the link prediction posterior probability of accurately predicting all interactions in the dev set:
-(1) $\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}})=\mathcal{E}_{\text{dev}})$
+$\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}})=\mathcal{E}_{\text{dev}})$ (1)
 where $\{(u,i)\}_{\text{dev}}$ is the dev set input user-item pairs, $\mathcal{G}_{\text{train}}$ is the graph used to train the model, $\mathcal{E}_{\text{dev}}$ is the ground truth positive edges in the dev set inputs.
 The formulation above means that the model maximizes the probability of accurately predicting edges on the dev set, while observing certain training graph $\mathcal{G}_{\text{train}}$.
-3.2. Recommendation with Data Augmentation
 We study a simple and effective approach to augment the graph, which adds new nodes and edges to augment the original graph $\mathcal{G}$ into a larger graph $\mathcal{G}\cup\mathcal{G^{+}}$.
 The training objective then becomes:
-(2) $\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}}\cup\mathcal{G}^{+})=\mathcal{E}_{\text{dev}})$
+$\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}}\cup\mathcal{G}^{+})=\mathcal{E}_{\text{dev}})$ (2)
 The augmented graph $\mathcal{G}^{+}$ is comprised of a unified set of intent nodes and heterogeneous connections to existing graph nodes, which is visualized in Figure 2 and to be described in Section 4.1.
-Figure 2.
-IKGR’s graph augmentation steps to build intent knowledge graph with LLM.
-IKGR overcomes knowledge gap via in-context learning, avoids synthetic noise by focusing on the simple node-level intent entity retrieval task, and being light weight and tuning-free.
-3.3. Knowledge Graph Convolution Layer
+3.2 Knowledge Graph Convolution Layer
 A Knowledge Graph (KG) is a directed graph composed of subject-property-object triple facts. Each triplet $(e_{h},e_{t},r)$ denotes a relationship $r$ from head entity $e_{h}$ to tail entity $e_{t}$.
-Similar to KGCN (Wang et al., 2019), we employ a knowledge graph convolution layer to capture structural proximity among entities in a knowledge graph.
-The model learns node embeddings $\bm{E}\in R^{T_{R}\times d}$ and relation embeddings $\bm{R}\in R^{N\times d}$, where $T_{R}$ is number of relation types, $N$ is number of nodes, and $d$ is the embedding dimensionality.
+Similar to KGCN [27], we employ a knowledge graph convolution layer to capture structural proximity among entities in a knowledge graph.
+The model learns node embeddings $\bm{E}\in R^{N\times d}$ and relation embeddings $\bm{R}\in R^{T_{R}\times d}$, where $T_{R}$ is number of relation types, $N$ is number of nodes, and $d$ is the embedding dimensionality.
 Denote the input features for some node $v$ as $\mathbf{v}$, the set of entities directly connected to $v$ as $\mathcal{S}_{v}$, then the output embedding is the summation-aggregated neighborhood entity and relation embeddings:
-(3) $\mathbf{v}^{\text{out}}=\sigma\left({\bf W}\cdot[{\bf v}+\text{softmax}(\bm{R}[\mathcal{S}_{v}]\bm{E}[\mathcal{S}_{v}]^{T})\bm{E}[\mathcal{S}_{v}]]+{\bf b}\right)$
+$\mathbf{v}^{\text{out}}=\sigma\left({\bf W}\cdot[{\bf v}+\text{softmax}(\bm{R}[\mathcal{S}_{v}]\bm{E}[\mathcal{S}_{v}]^{T})\bm{E}[\mathcal{S}_{v}]]+{\bf b}\right)$ (3)
 where $\bm{R}[\mathcal{S}_{v}]$ and $\bm{E}[\mathcal{S}_{v}]$ are the relationship embedding vector set and entity embedding vector set, each containing $|\mathcal{S}_{v}|$ of $d$-dimensional vectors.
 $\bf W$ and $\bf b$ are transformation weights and a bias term, respectively, and $\sigma$ is the activation function.
-3.4. Knowledge Graph Embedding Translation
-Due to the incomplete nature of KGs, KG completion is often leveraged as a self-supervised learning task, which predicts the missing entity $e_{h}$ or $e_{t}$ for a triplet $(e_{h},e_{t},r)$.
-To this end, TransE (Bordes et al., 2013), a popular knowledge graph embedding model commonly used for KG completion, enforces a translation in the embedding space: $\mathbf{e}_{h}+\mathbf{e}_{r}\approx\mathbf{e}_{t}$.
-This embedding translation is achieved via the following training objective: $\min\mathop{\Sigma}\limits_{\mathbf{e}_{h},\mathbf{r},\mathbf{e}_{t}}\parallel\mathbf{e}_{h}+\mathbf{r}-\mathbf{e}_{t}\parallel$.
-One issue with TransE embedding is that a single relation type may correspond to multiple head entities or tail entities, leading to significant 1-to-N, N-to-1, and N-to-N issues (Wang et al., 2014).
-As an improvement, TransH (Wang et al., 2014) learns different representations for an entity conditioned on different relations.
-It assumes that each relation owns a hyperplane, and the translation between head entity and tail entity is valid only if they are projected to the same hyperplane.
-It defines an energy score function for a triplet as follows:
-(4) $f(\mathbf{e}_{h},\mathbf{e}_{t},\mathbf{r})=\parallel\mathbf{e}_{h}^{\bot}+\mathbf{r}-\mathbf{e}_{t}^{\bot}\parallel$
-where a lower score of $f(e_{h},e_{t},r)$ indicates that the triplet is likely valid. $\mathbf{e}_{h}^{\bot}$ and $\mathbf{e}_{t}^{\bot}$ are projected entity vectors:
-(5) $\mathbf{e}_{h}^{\bot}=\mathbf{e}_{h}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{h}\mathbf{w}_{r}$
-(6) $\mathbf{e}_{t}^{\bot}=\mathbf{e}_{t}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{t}\mathbf{w}_{r}$
-where $\mathbf{w}_{r}$ and $\mathbf{r}$ are two learned vectors for a specified relation $r$.
-$\mathbf{w}_{r}$ denotes the projection vector of its corresponding hyperplane where $\mathbf{r}$ is the relationship embedding vector.
-Finally, the training of TransH encourages the discrimination between valid triplets and incorrect ones using margin-based ranking loss. Similar to KTUP (Cao et al., 2019), we employ:
-(7) $\mathcal{L}_{KG}=-\sum_{\begin{subarray}{c}(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})\in\mathcal{KG}\\ (\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})\in\mathcal{KG}^{-}\end{subarray}}\log\sigma[f(\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})-f(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})]$
-where $\mathcal{KG}^{-}$ contains corrupted triplets constructed by randomly sampling a tail entity and relation.
-In practice, weight decay and normalization-enforcing losses are also applied to prevent overfitting.
-4. Methodology
-In this section, we address the following key questions regarding LLM-based knowledge graph augmentation: Q1: How to enable node-level intent entity extraction with RAG?
-Q2: How to inject the learned intents as structural priors into KG modeling?
-Q3: How to improve scoring with intent awareness at test time?
-4.1. Intents extraction with RAG (Q1)
+4 Methodology
+The key components of IKGR involve grounded user/item intent extraction with RAG, KG densification and a GNN prediction layer.
+4.1 Intents Extraction with RAG
 Despite significant progress in incorporating side information into recommendation systems, introducing low-quality side information may even undermine what little signal we can glean from sparse interactions.
@@
 In order to address challenges posed by sparse user behavior scenarios, we leverage the LLM to perform data augmentation.
-Unlike existing LLM recommendation methods that directly synthesize user-item interactions (Wei et al., 2024), we take a more conservative approach, tasking the LLM with a simpler, more reliable role to minimize noise in the generated outputs.
+Unlike existing LLM recommendation methods that directly synthesize user-item interactions [19], we take a more conservative approach, tasking the LLM with a simpler, more reliable role to minimize noise in the generated outputs.
 This approach leverages the LLM’s pre-trained common sense knowledge for accurate distillation.
 We note that a forward call to the LLM with appropriate instructions is able to extract an intent entity set.
-For example, for item node $i$, we denote the set of intent entities as $\Omega_{i}$, $\Omega_{i}=LLM($###Item Description: {Item i description} Return a list of entities mentioned in the Item Description that the user may have intents to interact with)))).
-However, there are two gaps of this naive application: (1) the existence of domain-specific knowledge may make it difficult for the LLM to fully understand the context; (2) the connectivity of the extracted intents could follow a very long-tail distribution, meaning that the majority of intent nodes might only link to few user/item nodes.
-To mitigate these challenges, we propose the following improvements over the naive forward call.
-4.2. In-context Learning
-To keep the resource budget lightweight, we curate an in-context learning process without fine-tuning.
-For the enterprise use-case, which differs from the public dataset use-case, user profiles and item descriptions often contain abbreviations and domain-specific concepts unknown to LLMs.
-To address this knowledge gap, we curate a knowledge base of key-value pairs, where keys are the domain-specific abbreviations/concepts, and values are their explanations.
-We identify any such terminology present in the user profile or item description and append their corresponding explanations to the prompt under a ###$\#\#\#Concept\ Explanation$ tag.
+For example, for item node $i$, we denote the set of intent entities as $\Omega_{i}$, $\Omega_{i}=LLM($###Item Description: {Item i description} Return a list of entities mentioned in the Item Description that the user may have intents to interact with)).
+4.2 Factual Knowledge Access for Grounded Extraction
+The intent extraction step assumes a high quality Item Description paragraph describing the intent.
+However, the Item Description can often be poorly-formulated in the datasets, leading to a gap in properly extracting intent.
+(i) In the case of enterprise search, the existence of private enterprise knowledge makes it difficult for the LLM to fully understand the context, and (ii) in the open sourced datasets case, only the item names are available without appropriate description - though details of these items are often available via online search or baked into the LLM, so that an agent with web access is able to augment it easily [28].
+We bridge this gap by feeding the intent extraction module with enterprise private knowledge or open-world knowledge.
+Specifically, for our adapted enterprise use-case, user profiles and item descriptions often contain abbreviations and domain-specific concepts unknown to LLMs.
+We curate a knowledge base of key-value pairs, where keys are the domain-specific abbreviations/concepts, and values are their explanations.
+We identify any such terminology present in the user profile or item description and append their corresponding explanations to the prompt under a ###ConceptExplanation\#\#\#Concept\ Explanation tag.
 This enriches the context for the LLM, enabling better understanding without the need for fine-tuning.
-4.3. Intent connectivity enrichment via RAG
-A two-round process extracts intents and builds a knowledge graph.
+For open source datasets, the item name is first expanded into a paragraph summary by an LLM agent before it feeds into the entity extraction module.
+4.3 Intent Connectivity Enrichment via RAG
+The connectivity of the extracted intents could follow a very long-tail distribution, meaning that the majority of intent nodes might only link to few user/item nodes.
+To mitigate these challenges, a two-round process is used to extract and densify the intents KG.
 In the first round, a simple prompt template is used to generate specific entities, linking users and items to intent nodes.
@@
 In the second round, each user and item node is connected to additional intent entities from the fixed pool of intents generated in the first round.
-We call these new connections ”related intents”.
+We call these new connections "related intents".
 During this second round of related intent selection, the existing intents for a given user/item node are excluded from the retrieval step.
 Denote the intent entities obtained after the first round extraction as $\bar{\Omega}$.
-Given a user profile or an item description $text$, the second round intent extraction prompt can be formulated as: $LLM($###Knowledge Context: … ### Options: $R(x,\mathcal{N}\backslash\Omega_{x},K)$ What are the intents mentioned in $x$ that are the most relevant?)))) , where $R(x,\bar{\Omega}\backslash\Omega_{x}^{\prime},K)$ retrieves a group of $K$ intents that are semantically similar to the input text $x$, yet have not been extracted during the first round output ($\Omega_{x}^{\prime}$).
+Given a user profile or an item description texttext, the second round intent extraction prompt can be formulated as: LLM(LLM(###Knowledge Context: … ### Options: $R(x,\mathcal{N}\backslash\Omega_{x},K)$ What are the intents mentioned in $x$ that are the most relevant?)) , where $R(x,\bar{\Omega}\backslash\Omega_{x}^{\prime},K)$ retrieves a group of $K$ intents that are semantically similar to the input text $x$, yet have not been extracted during the first round output ($\Omega_{x}^{\prime}$).
+The full prompt is provided in Appendix A.
 New users and items undergo this two-round extraction to connect them to the existing, well-connected knowledge graph.
@@
 Both rounds prompt the LLM for structured output.
-Figure 3.
-The graph module of IKGR.
-After constructing the knowledge graph, the embeddings of user, item and intents are fed into the knowledge graph convolution layer followed by an embedding translation layer.
-4.4. Intent Translation GNN (Q2)
+4.4 Generating Recommendation Candidates with Graph Module
+After the KG has been built, a graph module is used to generate the recommendation candidate list.
+While multiple combinations of GNN architectures and loss functions could fit, we find a simple translation layer based architecture [29] that injects the learned intents as structural priors outperforms vanilla GNN options.
+We leverage this intent prior GNN as the default option in the experiment sections, and discuss the details in Appendix B.
+We also benchmark across three other options and show results in Section 5.3.
+5 Experiments
+In this section, we verify the performance of the proposed recommender using both proprietary and open-source data.
+For the proprietary data, we introduce IKGR to an enterprise knowledge search platform to test IKGR’s effectiveness in re-ranking search results.
+We also benchmark against existing baselines on four open-source recommendation datasets to verify the results in diverse scenarios.
+The details of these datasets are presented in Table 1.
+We use Llama-3.1-8B model for LLM inference, and all-mpnet-base-v2 from sentence transformer for encoding textual features.
+We retrieve the top 100 prebuilt intent candidates in the kNN retrieval step of RAG.
+We apply an 8:1:1 ratio when sampling the positive/negative edges for train/dev/test sets in the graph.
+[…]
+Datasets Search Beauty Books Steam Yelp2022
+#Users 36 03336\,033 40 22640\,226 251 394251\,394 281 428281\,428 1 987 8981\,987\,898
+#Items 872 678872\,678 54 54254\,542 25 60625\,606 13 04413\,044 150 347150\,347
+#Inter 3.5M 0.35M 3.2M 3.5M 6.9M
+Density 0.011% 0.02% 0.05% 0.095% 0.002%
+#Intents 495 285495\,285 39 30539\,305 45 93245\,932 53 94053\,940 65 04065\,040
+#IntEdges 6.9M 209K 390K 231K 409K
+AvgIntDeg 13.9 5.3 8.5 4.3 6.3
+5.1 Baselines and Datasets
+Enterprise Search.
+In the enterprise search setting, a set of developers query the search engine to search for datasets published by other developers within the enterprise.
+The search engine first uses BM25 [30] to retrieve a list of item candidates, then the IKGR is applied to rerank the search results.
+Evaluations are done over user’s feedback on how high the item of interest could be ranked among the final reranked list.
+In this setting, the items are enterprise datasets, which consist of text features such as dataset name, description, column names, ID labels, etc.
+The users are all registered within the enterprise and their developer profiles could be collected through separate channels to offer a hint about their dataset consumption preference.
+The IKGR is trained over historical user-dataset consumption interactions collected from separate channels.
+This dataset is labeled as Search in dataset description table and result tables.
+Books, Beauty.
+These datasets are obtained from Amazon review [^1] http://jmcauley.ucsd.edu/data/amazon/ in [31], which contains a variety of categories.
+We utilize the Books and Beauty categories.
+We leverage the features of title, sales type, sales rank, categories, price, and brand.
+Steam [^2] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data.
+This is a dataset collected from Steam [32], a large online video game distribution platform.
+We leverage the item features of app name, genres, publisher, sentiment, specs, tags.
+Yelp2022333https://www.yelp.com/dataset.
+This is a popular dataset for business recommendation.
+Given the large size, we use the transaction records after January 31st, 2022.
+We treat the categories of businesses as attributes for items, and user compliment types as attributes for users.
+To evaluate the ranking list of all the models, we employ Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR).
+Considering we only have one ground truth item for each user, HR@$k$ is equivalent to Recall@$k$ and proportional to Precision@$k$; MRR is equivalent to Mean Average Precision (MAP).
+We report HR and NDCG with k=1,5,10k={1,5,10}.
+For all these metrics, the higher the value, the better the performance.
+To verify the effectiveness of our method, we compare it with the following representative baselines.
+KGIN [13], which extracts non-readable intent in the latent embedding space.
+CSRec [10], which adds new nodes into KG with LLM, then post align the graph.
+HAKG [14], a hierarchy KG built upon the hyperbolic space.
+LLMRec [19], an LLM as interaction synthesizer approach.
+RippleNet [7], which uses a single layer of embedding translation by applying transH to user-item interaction.
+KGCN [27], KG based graph convolutional networks that captures inter-item relatedness effectively by mining their associated attributes on the KG.
+KTUP [29], which uses a single layer of embedding translation by applying TransH to user-item interaction.
+Table 2: Performance comparison of different methods. Bold scores are the best in each row, while underlined scores are the second best.
+Datasets Metric KGIN CSRec HAKG LLMRec RippleNet KGCN KTUP IKGR
+Search HR@1 0.0074 0.0079 0.0080 0.0075 0.0024 0.0051 0.0078 0.0086
+HR@5 0.0162 0.0156 0.0187 0.0158 0.0118 0.0161 0.0166 0.0202
+HR@10 0.0263 0.0258 0.0255 0.0253 0.0218 0.0232 0.0262 0.0267
+NDCG@5 0.0142 0.0144 0.0139 0.0148 0.0068 0.0101 0.0137 0.0151
+NDCG@10 0.0154 0.0153 0.0161 0.0164 0.0087 0.0123 0.0142 0.0172
+MRR 0.0135 0.0136 0.0128 0.0143 0.0091 0.0100 0.0130 0.0153
+Beauty HR@1 0.1103 0.1401 0.1623 0.1563 0.0532 0.0984 0.1783 0.1369
+HR@5 0.3092 0.2044 0.2984 0.2803 0.1972 0.3120 0.3388 0.3316
+HR@10 0.4194 0.4293 0.4817 0.4204 0.3695 0.4204 0.4610 0.4846
+NDCG@5 0.2398 0.2390 0.2184 0.2293 0.1307 0.1729 0.2583 0.2806
+NDCG@10 0.2643 0.2433 0.2580 0.2930 0.1713 0.2345 0.2814 0.2939
+MRR 0.2294 0.2300 0.2402 0.2203 0.1382 0.2254 0.2581 0.2641
+Books HR@1 0.1218 0.1194 0.1177 0.1020 0.0480 0.0853 0.1125 0.1251
+HR@5 0.2983 0.3093 0.2764 0.2674 0.1691 0.2465 0.2652 0.3197
+HR@10 0.3204 0.3449 0.3781 0.4094 0.3553 0.3582 0.3614 0.4248
+NDCG@5 0.1910 0.1980 0.1874 0.1877 0.1031 0.1579 0.1921 0.2097
+NDCG@10 0.2573 0.2673 0.2963 0.2673 0.1607 0.2016 0.2415 0.2814
+MRR 0.2130 0.2203 0.2599 0.2563 0.1407 0.1832 0.2193 0.2672
+Steam HR@1 0.0783 0.0847 0.0960 0.0744 0.0304 0.0641 0.1060 0.1095
+HR@5 0.2653 0.2483 0.2665 0.2174 0.1429 0.1966 0.2445 0.2759
+HR@10 0.2901 0.2899 0.3170 0.3237 0.2735 0.3218 0.3126 0.3574
+NDCG@5 0.1691 0.1335 0.1221 0.1694 0.0915 0.1496 0.1614 0.1735
+NDCG@10 0.1739 0.1562 0.1771 0.1884 0.1161 0.1905 0.1809 0.2212
+MRR 0.2007 0.1965 0.1882 0.1872 0.1294 0.1528 0.1957 0.2168
+Yelp2022 HR@1 0.0771 0.0936 0.1077 0.0724 0.0495 0.0917 0.1005 0.0989
+HR@5 0.2766 0.2355 0.2687 0.2211 0.1513 0.2405 0.2395 0.2869
+HR@10 0.3108 0.3362 0.3760 0.3228 0.2996 0.3395 0.3317 0.3966
+NDCG@5 0.1733 0.1823 0.1995 0.1931 0.1201 0.1397 0.2067 0.2078
+NDCG@10 0.2034 0.2164 0.2234 0.2127 0.1225 0.1861 0.2145 0.2277
+MRR 0.2093 0.2029 0.1842 0.1980 0.1356 0.1578 0.1906 0.2100
+5.2 Result Analysis
+Observing the results in Table 2, both datasets and models show statistical performance differences.
+From the dataset perspective, different datasets have different levels of knowledge gaps, with the Search dataset possibly having the largest knowledge gap for the LLM.
+This could be indicated by the number of intents and average intent node degrees in Table 1, where the Search dataset’s intent count is about one magnitude larger than other datasets.
+On the other hand, this is also observed by the performance difference of the IKGR model, where the same architecture is applied on different LLM augmented datasets.
+As IKGR performances are lower on the Search dataset, it hints that interaction prediction under the knowledge gap is more difficult.
+From the approaches perspective, IKGR shows new state-of-the-art performances especially on the Search dataset with the largest knowledge gap.
+On the other hand, across these datasets, IKGR performs better than implicit intent approach (KGIN) or intent unaware LLM approaches (LLMRec), demonstrating the gain from explicit entity node augmentation.
+5.3 Ablation Studies
+To investigate the influence of each component over the enterprise search data, we conduct a list of experiments that drop each component in the proposed pipeline and compare with the original pipeline on the Search and Beauty dataset.
+Results are presented in Table 3.
+Table 3: Ablation Study
+Dataset Versions HR@1 HR@5 HR@10 NDCG@5 NDCG@10 MRR
+Search Opt1. Int. Prior GNN 0.0086 0.0202 0.0267 0.0151 0.0172 0.0153
+Opt2. Vanilla GNN 0.0078 0.0184 0.048 0.0132 0.0152 0.0130
+Opt3. Vanilla Trans. 0.0082 0.0193 0.0256 0.0140 0.0164 0.0139
+Opt4. Vanilla Scoring 0.0084 0.0199 0.0258 0.0142 0.0161 0.0137
+No Related Intent 0.0077 0.0185 0.0254 0.0147 0.0163 0.0134
+No Intent 0.0073 0.0175 0.0240 0.0131 0.0150 0.0125
+Beauty Opt1. Int. Prior GNN 0.1369 0.3316 0.4846 0.2806 0.2939 0.2641
+Opt2. Vanilla GNN 0.1294 0.3201 0.4299 0.2537 0.2674 0.2502
+Opt3. Vanilla Trans. 0.1311 0.3193 0.4272 0.2519 0.2566 0.2439
+Opt4. Vanilla Scoring 0.1332 0.3214 0.4249 0.2555 0.2579 0.2522
+No Related Intent 0.1266 0.3284 0.4462 0.2643 0.2741 0.2536
+No Intent 0.1183 0.2984 0.4093 0.2463 0.2453 0.2399
+In Table 3, Opt1.
+Int.
+Prior GNN means the full IKGR version without component dropping.
+Opt2.
+Vanilla GNN means to use a plain GNN to make predictions over the generated user-intent-item graph.
+Opt3.
+Vanilla Trans.
+means removing both GNN and intent-aware scoring function, and only use a plain graph translation layer.
+Opt4.
+Vanilla Scoring means removing the intent prior scoring.
+These four options correspond to four candidates in modeling the intent graph, all detailed in Appendix B.
+No Related Intent means dropping the second round of related intent retrieval using RAG, and only exact intent edges are presented in the graph, without the related intent edges.
+No Intent drops all intent nodes and simply use the GNN of IKGR to predict over user-item graph.
+As observed in the results shown in Table 3, the relative contributions of each components of IKGR can be estimated.
+The most significant observation is that the intent edges augmentation steps as a whole helps to densify the graph almost twice, as reflected by the #IntEdges and #Inter in Table 1, and the MRR performance correspondingly improved 22% (0.0125 to 0.0153 in Search dataset).
+This confirms that adding intent relations to the graph improves the IKGR performance while densifying the graph with meaningful knowledge connectivities.
+The proposed intent node embedding improves the performance by offering a straightforward embedding structure between user and item vectors, hence guiding the learning procedure, as shown in Table 3 that learning the relation vectors from scratch and removing the intent based scoring both harms the performance.
+5.4 Cold Start Metrics
+We use the books dataset to evaluate the cold start setting.
+The evaluation set is chosen to be a subset of edges whose end nodes both have less than or equal to 3 interactions (node degrees).
+The results are presented in Table 4.
+As can be seen from the table, IKGR achieves better performance than other methods on the tail edge set, validating its effectiveness of graph augmentation in dealing with cold start.
+Table 4: Performance comparison over tail edges
+Setting HR@10 NDCG@10 MRR
+IKGR 0.4085 0.2791 0.2630
+ConvNCF $0.2699$ $0.1373$ $0.1382$
+FPMC $0.2810$ $0.1427$ $0.1326$
+KTUP $0.3114$ $0.2044$ $0.1679$
+Table 5: Results on different hyperparameter configurations
+Configuration HR@10 NDCG@10 MRR
+$k=120$, #conv=1=1 0.4262 0.2835 0.2711
+$k=100$, #conv=1=1 0.4248 0.2814 0.2672
+$k=80$, #conv=1=1 0.3610 0.2619 0.2520
+$k=50$, #conv=1=1 0.3023 0.2205 0.1945
+$k=100$, #conv=2=2 0.3735 0.2517 0.2482
+5.5 Hyperparameter Sensitiveness
+To test how model behaves under different hyperparameters, we computed top-k in the kNN retrieval step of RAG, and number of GNN layers in the model architecture.
+These experiments are conducted on the Books dataset.
+The hyperparameter sensitivity results are shown in Table 5.
+As seen from the table, the model kNN saturates at k=100, which validates our architecture choice of $k$ that both ensures performance and avoid over lengthy token sequences.
+5.6 Impact of Knowledge Base
+Table 6: Statistics of extracted intents
+Metric Value
+Average num entity with KB $15.5$
+Average num entity without KB $14.5$
+Avg Jaccard Similarity Coefficient $0.892$
+To quantify the influence of the knowledge base, we use the enterprise search data to compare the extracted intent sets for two scenarios: the knowledge base appended and dropped.
+We computed the average number of entity extracted, and the Jaccard Similarity Coefficient ( $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ ) for the entities under these two settings in Table 6.
+The result shows that the average number of intents does not differ much, and the Jaccard Similarity Coefficient is close to 1, meaning the knowledge base has limited impact on the extracted entity set.
+Our intention with the knowledge base is to serve as a lightweight information source in the event of heavy domain gap, rather than a bottleneck key component.
+Indeed, we have made the LLM generation task as a simple entity extraction task, hence the impact of missing structured knowledge is minimized.
+6 Conclusions
+In this work, we propose IKGR, a knowledge graph based recommender built with a Large Language Model (LLM).
+The proposed method features a data augmentation step to explicitly extract entities that the users have intents to interact with, and learns node embeddings over the knowledge graph using an embedding translation layer to combine the intent structure knowledge.
+This work takes the enterprise search personalization as a case study, and verifies that (1) when knowledge gap exists, using a simplified node-level augmentation task helps learn embeddings, while synthesizing interactions harms the model performance and introduces noise; (2) injecting intent structure prior into the modeling helps better capturing the semantic structure and boosts embedding learning.
+Appendix A Full Prompts And Visualizations
+For the enterprise search scenario, the full prompt in the first round of intent generation is provided as follows.
+### Task
+You are given a short paragraph of description of a dataset.
+Extract a few entities mentioned in the description that the users may be interested in when they consume this dataset.
+Acronyms can also be included in your answer, if any.
+Acronyms connected via underscores can be broken apart and partially used, so as to only include the core part into your answer.
+Please discard entities that have too general and broad meaning, and only pick those specific to this dataset.
+Please also discard entities that are over-specific, such as date or seemingly highly specified serial numbers/markers.
+Please return answer in the format of python list of string, such as: ["answer 1", "answer 2"]. Only return the python list of string. Do NOT return any other explaination words.
+### Dataset Description:
+{desc}
+### Answer:
+The full prompt in the second round of intent linkage densification is provided as follows.
+You are given a short paragraph of description of a dataset, and a list of entity candidates.
+Based on the available entity candidates, which of them are likely to attract the users attention, if the user has proven interest in the dataset?
+Please include related entities, which may not be diretly mentioned in the dataset description.
+Please make selections only from the available options given below.
+Please return answer in the format of python list of string.
+### Dataset Description:
+{desc}
+### Available Options:
+{RAG_options}
+### Important Notes:
+1. You can ONLY choose from the Available Options above. If certain entity is not mentioned in the Available Options, you CANNOT return that entity.
+2. Please return as many related entities as possible.
+3. Please only return the python list of string. Do NOT return any other explaination words.
+### Answer:
+For the open source dataset (e.g. Steam), the prompt for description summarization using the open world knowledge agent is as follows.
+{item_name} is a game on Steam. Collect necessary information for this game, and write one paragraph summary of why a user is potentially interested in it. Use objective tone.
+The above prompt converts the item name into a paragraph of LLM summary.
+Then LLM is used to extract the intent using a similar prompt as above.
+An example of extracted intents for steam dataset is as follows.
+Origin item name: Half-Life 2
+LLM summary: Half-Life 2 is a landmark first-person shooter that remains highly relevant today for its exceptional storytelling, physics-based gameplay, and rich world-building.
+Players could be interested in it because it offers more than just combat; the games silent protagonist, Gordon Freeman, navigates a deeply immersive narrative that unfolds entirely through gameplay, pulling you into a desperate human resistance against an alien empire.
+The revolutionary Gravity Gun allows for creative problem-solving and combat by manipulating objects, while the games detailed environments and memorable characters-like the resourceful Alyx Vance-create a compelling and lasting experience that set a new standard for the genre.
+Ultimately, its a masterclass in game design that is both historically significant and still incredibly fun to play.
+Extracted Intent entities: ["storytelling", "physics-based gameplay", "world-building", "combat", "narrative", "human resistance", "alien empire", "gravity gun", "problem-solving", "characters", "alyx vance", "game design"]
+Appendix B Intent Prior GNN
+Below we specify the details of intent prior GNN and how it improves scoring with intent awareness at test time.
+B.1 Knowledge Graph Embedding Translation Formulations
+Due to the incomplete nature of KGs, KG completion is often leveraged as a self-supervised learning task, which predicts the missing entity $e_{h}$ or $e_{t}$ for a triplet $(e_{h},e_{t},r)$.
+To this end, TransE [33], a popular knowledge graph embedding model commonly used for KG completion, enforces a translation in the embedding space: $\mathbf{e}_{h}+\mathbf{e}_{r}\approx\mathbf{e}_{t}$.
+This embedding translation is achieved via the following training objective: $\min\mathop{\Sigma}\limits_{\mathbf{e}_{h},\mathbf{r},\mathbf{e}_{t}}\parallel\mathbf{e}_{h}+\mathbf{r}-\mathbf{e}_{t}\parallel$.
+One issue with TransE embedding is that a single relation type may correspond to multiple head entities or tail entities, leading to significant 1-to-N, N-to-1, and N-to-N issues [34].
+As an improvement, TransH [34] learns different representations for an entity conditioned on different relations.
+It assumes that each relation owns a hyperplane, and the translation between head entity and tail entity is valid only if they are projected to the same hyperplane.
+It defines an energy score function for a triplet as follows:
+$f(\mathbf{e}_{h},\mathbf{e}_{t},\mathbf{r})=\parallel\mathbf{e}_{h}^{\bot}+\mathbf{r}-\mathbf{e}_{t}^{\bot}\parallel$ (4)
+where a lower score of $f(e_{h},e_{t},r)$ indicates that the triplet is likely valid. $\mathbf{e}_{h}^{\bot}$ and $\mathbf{e}_{t}^{\bot}$ are projected entity vectors:
+$\mathbf{e}_{h}^{\bot}=\mathbf{e}_{h}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{h}\mathbf{w}_{r}$ (5)
+$\mathbf{e}_{t}^{\bot}=\mathbf{e}_{t}-\mathbf{w}_{r}^{\mathrm{T}}\mathbf{e}_{t}\mathbf{w}_{r}$ (6)
+where $\mathbf{w}_{r}$ and $\mathbf{r}$ are two learned vectors for a specified relation $r$.
+$\mathbf{w}_{r}$ denotes the projection vector of its corresponding hyperplane where $\mathbf{r}$ is the relationship embedding vector.
+Finally, the training of TransH encourages the discrimination between valid triplets and incorrect ones using margin-based ranking loss. Similar to KTUP [29], we employ:
+$\mathcal{L}_{KG}=-\sum_{\begin{subarray}{c}(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})\in\mathcal{KG}\\ (\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})\in\mathcal{KG}^{-}\end{subarray}}\log\sigma[f(\mathbf{e}_{h},\mathbf{e}^{-}_{t},\mathbf{r}^{-})-f(\mathbf{e}_{h},\mathbf{e}^{+}_{t},\mathbf{r}^{+})]$ (7)
+where $\mathcal{KG}^{-}$ contains corrupted triplets constructed by randomly sampling a tail entity and relation.
+In practice, weight decay and normalization-enforcing losses are also applied to prevent overfitting.
+Figure 3: The intent prior GNN in IKGR.
+This module is the knowledge graph convolution layer followed by an embedding translation layer, where the intent translation is encoded as structure prior in the embedding space.
+B.2 Intent Prior Injected Translation layers
 Having obtained the intent nodes from items and users, we train a knowledge graph-based GNN to predict user-item interaction probabilities.
@@
 Then, to enforce the relationship of more similar intent pairs inducing a smaller translation vector, the resulting vector $\mathbf{r}^{u,i}$ is computed via:
-(8) $\mathbf{r}^{u,i}=\text{softmax}(\mathbf{P}^{u,i})^{T}\mathbf{D}^{u,i}$
-4.5. Intent Aware Scoring (Q3)
+$\mathbf{r}^{u,i}=\text{softmax}(\mathbf{P}^{u,i})^{T}\mathbf{D}^{u,i}$ (8)
+B.3 Intent Aware Scoring
 To further leverage signal from the intents extracted during KG construction, we incorporate them into the user-item interaction scoring function, in conjunction with the more traditional embedding similarity score.
@@
 The embedding matching score between $u$ and $i$ is the cosine similarity between their embeddings:
-(9) $z_{u,i}=\frac{\mathbf{e}_{u}\cdot\mathbf{e}_{i}}{\parallel\mathbf{e}_{u}\parallel\cdot\parallel\mathbf{e}_{i}\parallel}$
-We denote a single intent extracted from $q$ as $n^{(q)}$, and the collection of all intents for $q$ as $\Omega_{q}$.
-Similarly, a single intent extracted from $u$ and their collections is denoted as $n^{(u)}$ and $\Omega_{u}$, and a single intent extracted from $i$ and their collections is denoted as $n^{(i)}$, and $\Omega_{i}$.
-Additionally, $\mathbf{e}_{x}$ is used to represent the embedding of some entity $x$.
+$z_{u,i}=\frac{\mathbf{e}_{u}\cdot\mathbf{e}_{i}}{\parallel\mathbf{e}_{u}\parallel\cdot\parallel\mathbf{e}_{i}\parallel}$ (9)
+We denote a single intent extracted from item $i$ as $n^{(i)}$, and the collection of all intents for $i$ as $\Omega_{i}$.
+Similarly, a single intent extracted from $u$ and their collections is denoted as $n^{(u)}$ and $\Omega_{u}$.
+Additionally, $\mathbf{e}_{n}$ is used to represent the embedding of some entity $n$.
 The intent matching score between $u$ and $i$ is determined by:
-(10) $y_{u,i}=\mathop{max}\limits_{n^{(u)}\in\Omega_{u},n^{(i)}\in\Omega_{i}}\frac{\mathbf{e}_{n^{(u)}}\cdot\mathbf{e}_{n^{(i)}}}{||\mathbf{e}_{n^{(u)}}\parallel\cdot\parallel\mathbf{e}_{n^{(i)}}\parallel}\cdot 0.5^{\mathbb{I}(\Omega_{u}\cap\Omega_{i}=\varnothing)}$
+$y_{u,i}=\mathop{max}\limits_{n^{(u)}\in\Omega_{u},n^{(i)}\in\Omega_{i}}\frac{\mathbf{e}_{n^{(u)}}\cdot\mathbf{e}_{n^{(i)}}}{||\mathbf{e}_{n^{(u)}}\parallel\cdot\parallel\mathbf{e}_{n^{(i)}}\parallel}\cdot 0.5^{\mathbb{I}(\Omega_{u}\cap\Omega_{i}=\varnothing)}$ (10)
+Equation 10 has two components, the similarity component and the non-overlap punishment component.
+In the event that $u$ and $i$ share a single intent (no need to have completely equal intent set), the non-overlap punishment component will be disabled (=1), otherwise if no intents are shared, it will punish by 0.5.
+The choice of 0.5 is decided based on empirical comparison along the development of the approach, which shows priority over too severe punishment values (e.g.
+0.1).
 Finally, the score for the triplet $(u,i)$ is the hybrid mixture of the embedding score and intent matching score:
-(11) $score(u,i)=y_{u,i}+\lambda z_{u,i}$
-where $\lambda$ is an empirical mixer coefficient.
-4.6. Inference Procedure And Complexity Analysis
-5. Experiments
-In this section, we verify the performance of the proposed recommender using both proprietary and open-source data.
-For the proprietary data, we introduce IKGR to an enterprise knowledge search platform to test IKGR’s effectiveness in re-ranking search results.
-We also benchmark against existing baselines on four open-source recommendation datasets to verify the results in diverse scenarios.
-The details of these datasets are presented in Table 1.
-We use Llama-3.1-8B model for LLM inference, and all-mpnet-base-v2 from sentence transformer for encoding textual features.
-We retrieve the top 100 prebuilt intent candidates in the kNN retrieval step of RAG.
-We apply an 8:1:1 ratio when sampling the positive/negative edges for train/dev/test sets in the graph.
-The coefficient $\lambda$ in Equation 11 is set to 0.1 based on grid search result on the validation set.
-[…]
-Datasets Search Beauty Books Steam Yelp2022
-#Users 36 0333603336\,03336 033 40 2264022640\,22640 226 251 394251394251\,394251 394 281 428281428281\,428281 428 1 987 89819878981\,987\,8981 987 898
-#Items 872 678872678872\,678872 678 54 5425454254\,54254 542 25 6062560625\,60625 606 13 0441304413\,04413 044 150 347150347150\,347150 347
-#Inter 3.5M 0.35M 3.2M 3.5M 6.9M
-Density 0.011% 0.02% 0.05% 0.095% 0.002%
-#Intents 495 285495285495\,285495 285 39 3053930539\,30539 305 45 9324593245\,93245 932 53 9405394053\,94053 940 65 0406504065\,04065 040
-#IntEdges 6.9M 209K 390K 231K 409K
-AvgIntDeg 13.9 5.3 8.5 4.3 6.3
-5.1. Baselines and Datasets
-Enterprise Search.
-In the enterprise search setting, a set of developers query the search engine to search for datasets published by other developers within the enterprise.
-The search engine first uses BM25 (Whissell and Clarke, 2011) to retrieve a list of item candidates, then the IKGR is applied to rerank the search results.
-Evaluations are done over user’s feedback on how high the item of interest could be ranked among the final reranked list.
-In this setting, the items are enterprise datasets, which consist of text features such as dataset name, description, column names, ID labels, etc.
-The users are all registered within the enterprise and their developer profiles could be collected through separate channels to offer a hint about their dataset consumption preference.
-The IKGR is trained over historical user-dataset consumption interactions collected from separate channels.
-This dataset is labeled as Search in dataset description table and result tables.
-Books, Beauty.
-These datasets are obtained from Amazon review [^1] http://jmcauley.ucsd.edu/data/amazon/ in (McAuley et al., 2015), which contains a variety of categories.
-We utilize the Books and Beauty categories.
-We leverage the features of title, sales type, sales rank, categories, price, and brand.
-Steam [^2] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data.
-This is a dataset collected from Steam (Kang and McAuley, 2018), a large online video game distribution platform.
-We leverage the item features of app name, genres, publisher, sentiment, specs, tags.
-Yelp2022333https://www.yelp.com/dataset.
-This is a popular dataset for business recommendation.
-Given the large size, we use the transaction records after January 31st, 2022.
-We treat the categories of businesses as attributes for items, and user compliment types as attributes for users.
-To evaluate the ranking list of all the models, we employ Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR).
-Considering we only have one ground truth item for each user, HR@$k$ is equivalent to Recall@$k$ and proportional to Precision@$k$; MRR is equivalent to Mean Average Precision (MAP).
-We report HR and NDCG with $k={1,5,10}$.
-For all these metrics, the higher the value, the better the performance.
-To verify the effectiveness of our method, we compare it with the following representative baselines:
-POP: It is the simplest baseline that ranks items according to their popularity judged by the number of interactions.
-BPR-MF (Rendle et al., 2012): It optimizes the matrix factorization with implicit feedback using a pairwise ranking loss.
-FM (Rendle, 2010): The factorization machine models all interactions between variables using factorized parameters.
-ConvNCF (He et al., 2018): It employs a convolutional neural network to learn high-order correlations among embedding dimensions neural collaborative filtering.
-FPMC (Rendle et al., 2010): It captures users’ general taste as well as their sequential behaviors by combing MF with first-order Markov chains.
-RippleNet (Wang et al., 2018): It uses a single layer of embedding translation by applying transH to user-item interaction.
-KGCN (Wang et al., 2019): The KG based graph convolutional networks captures inter-item relatedness effectively by mining their associated attributes on the KG.
-KTUP (Cao et al., 2019): It uses a single layer of embedding translation by applying TransH to user-item interaction.
-Table 2. Performance comparison of different methods. Bold scores are the best in each row, while underlined scores are the second best.
-Datasets Metric POP BPR-MF ConvNCF FPMC RippleNet KGCN KTUP IKGR
-Search HR@1 0.0001 0.0041 0.0075 0.0048 0.0024 0.0051 0.0078 0.0086
-HR@5 0.0023 0.0100 0.0149 0.0125 0.0118 0.0161 0.0166 0.0202
-HR@10 0.0076 0.0174 0.0262 0.0225 0.0218 0.0232 0.0262 0.0267
-NDCG@5 0.0029 0.0068 0.0090 0.0074 0.0068 0.0101 0.0137 0.0151
-NDCG@10 0.0046 0.0094 0.0112 0.0107 0.0087 0.0123 0.0142 0.0172
-MRR 0.0046 0.0108 0.0070 0.0097 0.0091 0.0100 0.0130 0.0153
-Beauty HR@1 0.0405 0.0552 0.1005 0.1009 0.0532 0.0984 0.1783 0.1369
-HR@5 0.0377 0.1844 0.2754 0.2379 0.1972 0.3120 0.3388 0.3316
-HR@10 0.1074 0.3540 0.4387 0.3957 0.3695 0.4204 0.4610 0.4846
-NDCG@5 0.0763 0.1245 0.1758 0.1335 0.1307 0.1729 0.2583 0.2806
-NDCG@10 0.0443 0.1731 0.1961 0.2123 0.1713 0.2345 0.2814 0.2939
-MRR 0.0909 0.1792 0.1736 0.1752 0.1382 0.2254 0.2581 0.2641
-Books HR@1 0.0015 0.0324 0.1124 0.0620 0.0480 0.0853 0.1125 0.1251
-HR@5 0.0524 0.1663 0.2248 0.1859 0.1691 0.2465 0.2652 0.3197
-HR@10 0.1073 0.2649 0.3181 0.3217 0.3553 0.3582 0.3614 0.4248
-NDCG@5 0.0314 0.0983 0.1371 0.1268 0.1031 0.1579 0.1921 0.2097
-NDCG@10 0.0532 0.1603 0.1633 0.1703 0.1607 0.2016 0.2415 0.2814
-MRR 0.0788 0.1317 0.1567 0.1532 0.1407 0.1832 0.2193 0.2672
-Steam HR@1 0.0091 0.0547 0.0860 0.0526 0.0304 0.0641 0.1060 0.1095
-HR@5 0.0370 0.1196 0.1976 0.1698 0.1429 0.1966 0.2445 0.2759
-HR@10 0.0901 0.2379 0.3374 0.2717 0.2735 0.3218 0.3126 0.3574
-NDCG@5 0.0291 0.1035 0.1121 0.1094 0.0915 0.1496 0.1614 0.1735
-NDCG@10 0.0459 0.1244 0.1471 0.1384 0.1161 0.1905 0.1809 0.2212
-MRR 0.0507 0.1265 0.1382 0.1272 0.1294 0.1528 0.1957 0.2168
-Yelp2022 HR@1 0.0171 0.0636 0.1177 0.0524 0.0495 0.0917 0.1005 0.0989
-HR@5 0.0366 0.1555 0.1987 0.1911 0.1513 0.2405 0.2395 0.2869
-HR@10 0.1018 0.2462 0.3576 0.3028 0.2996 0.3395 0.3317 0.3966
-NDCG@5 0.0233 0.0823 0.1395 0.1131 0.1201 0.1397 0.2067 0.2078
-NDCG@10 0.0334 0.1364 0.1434 0.1527 0.1225 0.1861 0.2145 0.2277
-MRR 0.0593 0.1327 0.1542 0.1200 0.1356 0.1578 0.1906 0.2100
-Observing the results in Table 2, both datasets and models show statistical performance differences.
-From the dataset perspective, different datasets have different levels of knowledge gaps, with the Search dataset possibly having the largest knowledge gap for the LLM.
-This could be indicated by the number of intents and average intent node degrees in Table 1, where the Search dataset’s intent count is about one magnitude larger than other datasets.
-On the other hand, this is also observed by the performance difference of the IKGR model, where the same architecture is applied on different LLM augmented datasets.
-As IKGR performances are lower on the Search dataset, it hints that interaction prediction under the knowledge gap is more difficult.
-From the approaches perspective, IKGR shows new state-of-the-art performances especially on the Search dataset with the largest knowledge gap.
-On the other hand, across these datasets, the knowledge graph embedding translation approaches (IKGR, TKUP) or GNN based approaches (KGCN, RippleNet) performs better than factorization machines or collaborative filtering based approaches (BPR-MF, FPMC, ConvNCF), demonstrating the gain from structural embedding learning.
-5.2. Ablation Studies
-Table 3. Ablation Study
-Metrics HR@1 HR@5 HR@10 NDCG@5 NDCG@10 MRR
-IKGR 0.0086 0.0202 0.0267 0.0151 0.0172 0.0153
-Exact Int. Only 0.0077 0.0185 0.0254 0.0147 0.0163 0.0134
-No Aug. 0.0073 0.0175 0.0240 0.0131 0.0150 0.0125
-Inter. Aug. 0.0070 0.0166 0.0220 0.0126 0.0141 0.0117
-No Int. scoring 0.0082 0.0193 0.0256 0.0140 0.0164 0.0139
-No Int. Trans. 0.0078 0.0184 0.048 0.0132 0.0152 0.0130
-To investigate the influence of each component over the enterprise search data, we conduct a list of experiments that drop each component in the proposed pipeline and compare with the original pipeline on the Search dataset.
-Results are presented in Table 3.
-In Table 3, IKGR means the original approach without component dropping, Exact Int.
-Only means dropping the second round of related intent retrieval using RAG, and only exact intent edges are presented in the graph, without the related intent edges.
-No Aug.
-means training the architecture of IKGR, which is a naive knowledge graph convolution followed by an embedding translation layer, without adding the intent node in the graph.
-Inter.
-Aug.
-means to employ LLM only as a user-item implicit feedback augmentor without adding the intent nodes (Wei et al., 2024).
-In this setting the LLM is asked to select items that the user might be likely or unlikely to interact with.
-No Int.
-scoring means using a simple cosine similarity based score and we do not apply intent-aware scoring function as in Equation 11.
-No Int.
-Trans.
-means intent translation is not leveraged to compute the translation as in Equation 8.
-As observed in the results shown in Table 3, the relative contributions of each components of IKGR can be estimated.
-The most significant observation is that the intent edges augmentation steps as a whole helps to densify the graph almost twice, as reflected by the #IntEdges and #Inter in Table 1, and the MRR performance correspondingly improved 22% (0.0125 to 0.0153) shown in Table 3.
-This confirms that adding intent relations to the graph improves the IKGR performance while densifying the graph with meaningful knowledge connectivities.
-The proposed intent node embedding improves the performance by offering a straightforward embedding structure between user and item vectors, hence guiding the learning procedure, as shown in Table 3 that learning the relation vectors from scratch and removing the intent based scoring both harms the performance.
-6. Conclusions
-In this work, we propose IKGR, a knowledge graph based recommender built with a Large Language Model (LLM).
-The proposed method features a data augmentation step to explicitly extract entities that the users have intents to interact with, and learns node embeddings over the knowledge graph using an embedding translation layer to combine the intent structure knowledge.
-This work takes the enterprise search personalization as a case study, and verifies that (1) when knowledge gap exists, using a simplified node-level augmentation task helps learn embeddings, while synthesizing interactions harms the model performance and introduces noise; (2) injecting intent structure prior into the modeling helps better capturing the semantic structure and boosts embedding learning.
+$score(u,i)=y_{u,i}+\lambda z_{u,i}$ (11)
+where $\lambda$ is an empirical mixer coefficient. In the experiments, the coefficient $\lambda$ is set to 0.1 based on grid search result on the validation set.
+In the ablation study section Section 5.3, Opt2.
+Vanilla GNN means translation layers, simply use a two-layer GNN trained with standard ranking loss.
+Opt3.
+Vanilla Trans.
+means simply use a translation layer without GNN and the scoring function of Equation 11.
+Opt4.
+Vanilla Scoring means keep both GNN and translation layer, but removing intent-aware scoring function of Equation 11.
+Appendix C Related Works Continued
+Recommender Approaches Addressing Data Sparsity.
+Recent works tackle the data sparsity issue through enhanced model architectures, graph-based techniques, and knowledge-aware methods.
+The Item History Model (IHM) improves cold-start item recommendations by directly injecting user-interaction data into the item tower and employing an inductive structure for real-time inference [6].
+Cold Brew addresses sparsity in graph neural networks by distilling node representations, mitigating the impact of missing or noisy neighbors [35].
+RippleNet further alleviates sparsity by propagating user preferences through knowledge graph relations, enriching item representations beyond collaborative signals [7].
+These approaches demonstrate that leveraging historical interactions, distilling structural knowledge, and integrating external information effectively mitigate data sparsity in recommendation systems.
+Directly Leveraging LLM as the Recommender.
+A-LLMRec enhances LLM-based recommendation by leveraging embeddings from state-of-the-art collaborative filtering models, excelling in both cold and warm scenarios while maintaining efficiency and model-agnostic integration [36].
+CoRAL introduces collaborative retrieval-augmented prompting, addressing LLMs’ reliance on semantic information by incorporating user-item interactions through reinforcement learning-based retrieval policies, significantly improving long-tail recommendation [37].
+TALLRec proposes a tuning framework to align LLMs with recommendation-specific tasks, demonstrating strong generalization and efficiency even with limited training data [38].
+These approaches collectively highlight the potential of LLMs as standalone recommenders by addressing cold-start challenges, enhancing collaborative reasoning, and improving alignment with recommendation-specific objectives.
+Personalized Re-ranking Recommender System.
+Re-ranking in recommender systems aims to refine an initially ranked list to better capture user preferences and item relationships.
+Traditional ranking methods optimize global performance but often overlook the mutual influence between items and user-specific intent [4].
+Recent approaches address these limitations by integrating semantic tag information [39], multi-stage ranking optimization [40], and item relationships [5].
+Transformer-based models effectively model global item interactions, while graph-based methods leverage item relationships for improved ranking [4, 5].
+Furthermore, self-supervised learning enhances sequential recommendation by mitigating data sparsity issues [41].
+These methods collectively demonstrate that incorporating user personalization, item dependencies, and efficient ranking strategies significantly enhances re-ranking effectiveness.